### v1

# Towards Cross-Lingual LLM Evaluation for European Languages

Abstract
The rise of Large Language Models (LLMs) has revolutionized natural language processing across numerous languages and tasks. However, evaluating LLM performance in a consistent and meaningful way across multiple European languages remains challenging, especially due to the scarcity of multilingual benchmarks. We introduce a cross-lingual evaluation approach tailored for European languages. We employ translated versions of five widely-used benchmarks to assess the capabilities of 40 LLMs across 21 European languages. Our contributions include examining the effectiveness of translated benchmarks, assessing the impact of different translation services, and offering a multilingual evaluation framework for LLMs that includes newly created datasets: EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K. The benchmarks and results are made publicly available to encourage further research in multilingual LLM evaluation.
Towards Cross-Lingual LLM Evaluation for European Languages
Klaudia Thellmann1 Bernhard Stadler1 Michael Fromm2,3 Jasper Schulze Buschhoff2 Alex Jude2 Fabio Barth4 Johannes Leveling2 Nicolas Flores-Herr2 Joachim Köhler2 René Jäkel1 Mehdi Ali2,3 1TU Dresden, Germany, 2 Fraunhofer IAIS, Germany 3Lamarr Institute, Germany, 4DFKI, Germany thanks: †Equal contribution.
1 Introduction
The development of Large Language Models (LLMs) has brought transformative advancements to natural language processing, enabling sophisticated tasks such as question-answering, summarization, and machine translation. However, despite these advances, the evaluation of LLMs across languages, especially those beyond English and Chinese, remains a significant challenge. Initiatives like CLEF [^1] https://clef2024.imag.fr/, FIRE [^2] http://fire.irsi.res.in/, and WMT [^3] https://www2.statmt.org/wmt24/ have focused on multilingual evaluations, yet there is a lack of dedicated benchmarks for systematically comparing LLM performance across the languages of Europa.
Creating bespoke benchmarks for each language is a costly and time-consuming endeavor. This not only limits the scalability of evaluations but also results in a fragmented understanding of model performance across different languages. In many cases, evaluations are conducted in isolation, without a basis for meaningful comparison between languages. Moreover, human evaluation is widely recognized as the gold standard for assessing language model performance, as it captures nuances that automated benchmarks may miss. However, human evaluations are resource-intensive and difficult to scale across multiple languages.
To address these challenges, we explore whether automatically translated benchmarks can provide a viable method to estimate human preferences in chat scenarios. While human evaluation would likely be the most accurate way to assess LLM performance, the resource demands of manual dataset creation and evaluation make it impractical for large-scale multilingual comparisons. Instead, this approach aims to reduce these resource demands while ensuring comprehensive evaluation coverage across all languages.
In this paper, we focus on evaluating multilingual LLMs by utilizing translated versions of existing benchmarks. Our goal is to determine the efficacy of translated benchmarks and assess whether they can substitute manually generated ones, allowing for a more scalable and uniform evaluation across multiple languages.
We make the following contributions:
  • We evaluate the performance of 40 state-of-the-art LLMs for five multilingual datasets covering English and 20 additional languages in Section 4.
  • We assess whether translated benchmarks reflect human preferences to evaluate their viability in Section 5.
  • We examine how the choice of translation service (i.e., DeeplPro and ChatGPT) affects the performance of current LLMs in accurately responding to various benchmark tasks in Section 6.
  • We release EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K in 20 European languages as evaluation benchmarks for the NLP community [^4] https://huggingface.co/collections/openGPT-X/eu20-benchmarks-67093b13db8ff192dc39312d.
2 Related Work
Multilingual benchmarks have predominantly been generated using two widely adopted methodologies, besides various intermediate approaches: either by using human annotators of multiple languages to develop a multilingual dataset from scratch Kocmi et al. (2023); Goyal et al. (2022); Conneau et al. (2018) or by using LLMs to translate an existing benchmark into multiple languages Lai et al. (2023); Tiedemann (2012).
Creating a multilingual dataset with human annotators requires a significant amount of time, which impacts both the duration of production and the cost of the dataset Yang et al. (2019). In addition, homogeneous benchmarks are required in each language instead of heterogeneous multilingual benchmarks for a cross-lingual analysis Lewkowycz et al. (2022); Tiedemann (2012).
This method is increasingly inefficient because model scores can become saturated over time Ott et al. (2022), or benchmarks loose their relevance too fast if a desired performance is reached Lewkowycz et al. (2022). An alternative is the machine translation of existing benchmarks. This approach is both more cost-effective and time-efficient. However, the main disadvantage is the translation quality, which is highly dependent on the capability of the language model used Meng et al. (2022). In the past, some data sets have shown quality gaps in translation accuracy, with special consideration given to languages with medium and low resources Team et al. (2022).
Those languages have only a few to very few data sets available for pre-training and evaluation. They are much more difficult to be translated by an LLM, when taking cultural nuances of each language in consideration. Translating these languages, in particular, can lead to imbalanced evaluations across the benchmarks Team et al. (2022); Goyal et al. (2022); Conneau et al. (2018). Therefore, the selection of languages is a central decision to make when creating a multilingual benchmark.
In addition to the choice of languages, it is crucial to consider the range of tasks that need to be covered for a benchmark. For example, creating a multilingual benchmark for a specific task provides valuable insights into model performance for that exact task but says little about a model’s general applicability Conneau et al. (2018). This also includes benchmarks that, for example, cover the question-answering (QA) task for a wide range of domains, as only the QA task is tested here Lewis et al. (2020).
Another approach is to cover multiple tasks to evaluate language models broadly. An example is the OKAPI dataset published by Lai et al. (Lai et al., 2023). For this benchmark Three datasets (AI2 Reasoning Challenge (ARC), HellaSwag, and MMLU) were translated into 26 languages: eight high-resource, 11 medium-resource, and seven low-resource languages Lai et al. (2023).
Building on this approach, we have expanded the linguistic reach through additional translations with special consideration on European languages and increased the variety of tasks for a wider capability test of LLMs. Furthermore, we show that our data pre-processing has improved quality in certain language areas (See Section 3).
3 Experimental Setup
Firstly, we describe in Section 3.1 the translation process of the benchmarks, secondly Section 3.2 details our evaluation framework and process.
3.1 Translation Process
We chose DeepL as translation service for its balance between translation accuracy and scalability. We translated five well-known datasets, ARC Clark et al. (2018), HellaSwag  Zellers et al. (2019), TruthfulQA Lin et al. (2022), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2020) from English into 20 European languages. These datasets encompass a mix of multiple-choice and open-ended generation tasks, each of which presented unique translation challenges. The translations maintained the original structure of each task to ensure the benchmarks were as consistent as possible across languages. Using DeepL’s XML-tag handling, we avoided issues with prompt formatting and ensured that key contextual elements, such as ’Question’ and ’Response’ prompts, were properly localized (cf. Table 5).
To enhance the visibility and accessibility of our benchmarks, we launched a multilingual leaderboard [^5] https://huggingface.co/spaces/openGPT-X/european-llm-leaderboard on Hugging Face, inspired by the popular Open LLM leaderboard [^6] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard.
3.2 Evaluation Process
We conducted the model evaluations on the HPC Cluster Alpha Centauri (TUD/ZIH) [^7] https://compendium.hpc.tu-dresden.de, the JUWELS Booster (JSC) [^8] https://apps.fz-juelich.de/jsc/hps/juwels/index.html, and, for larger models, on the JURECA Cluster (WEST AI/JSC) [^9] https://westai.de/services/hardware/, using systems powered by NVIDIA A100 and H100 Tensor Core GPUs. For model evaluation, we used the LM Evaluation Harness Framework from Eleuther.AI [^10] https://github.com/OpenGPTX/lm-evaluation-harness in versions v0.4.1 and v0.4.3, evaluating in a multi-GPU setting using hf/accelerate with both data and model parallelism. For larger models, we deployed a local vllm-based inference server (v0.5.3) [^11] https://github.com/vllm-project/vllm/releases/tag/v0.5.5 using the LM Evaluation Harness local completions implementation (commit 42dc24/main) to perform pipeline-parallel evaluations across multiple GPUs on a single node.
4 European LLM Evaluation Results
This section provides an overview of the performance of large language models (LLMs) across various sizes, tasks and languages. Section 4.1 focuses on the evaluation across all 21 European languages and indicates which models provide a competitive performance in a multilingual setting. Section 4.2 shows the connection between the proportion of languages in the CommonCrawl dataset and the performance of the current best open weights models. Lastly, Section 4.3 highlights how the performance of the models depend on the language family. Given that many models provide little to no information about the composition of their training corpora, we focus our detailed analysis on the LLMs that demonstrate the most reliable performance across 21 languages.
4.1 Global European LLM Evaluation
This section provides an overview of the performance of several LLMs on various tasks and languages, grouped by their number of parameters in small-sized ($\leq$9B parameters), medium-sized ($>$9B & $<$27B parameters) and large ($>$27B parameters) models. The models are evaluated on our newly introduced tasks EU21-ARC, EU21-GSM8K, EU21-HellaSwag, EU21-MMLU, and EU21-TruthfulQA. An extensive overview of all models are provided in the Appendix as Tables (cf. LABEL:tab:acc_std_21lang_1, LABEL:tab:acc_std_21lang_2) and task specific heatmaps (cf. Figure 5, 6 7, 8, 9).
Figure 1: Accuracy results for four small sized LLMs on the EU21 tasks.
Small-Sized Models ($\leq$9B Parameters)
Among the small-sized models, the focus is on Gemma-2-9b-Instruct  Team et al. (2024b), Meta-Llama-3.1-8B-Instruct  Dubey et al. (2024), Mistral-NeMo-Minitron-8B-Base Sreenivas et al. (2024), and Qwen2-7B-Instruct Yang et al. (2024), further models with a lower average performance like Gemma-7b Team et al. (2024a) are in the appendix LABEL:tab:acc_std_21lang_1 and LABEL:tab:acc_std_21lang_2
Gemma-2-9b-Instruct consistently performs well (cf. Figure 1), by having a narrow interquartile ranges, signaling low variability and therefore stable performance in all languages. On the multilingual grade school math word benchmark EU20-GSM8K the median accuracy is worse than Meta-Llama-3.1-8B-Instruct, indicating lower math capabilities. In EU20-MMLU the model shows a robust behaviour, but the median is slightly lower than Qwen2-7B-Instruct. Both latter benchmarks indicate that the capacity of small models might not allow for both a reliable performance on all languages, aswell as the capability to handle specialized knowledge reliable.
Meta-Llama-3.1-8B-Instruct also delivers solid results, occasionally surpassing Gemma-2-9b-Instruct, such as in the EU21-GSM8K task where it achieves the highest median performance. While capable in math related tasks, the model shows shortcomings across all 21 languages compared to Gemma-2-9b.
Mistral-NeMo-Minitron-8B-Base performs competitively in certain tasks, particularly in more widely spoken languages, indicated by competitive maximum scores. However, its overall performance is weaker compared to other two models, with significant drops in less-resourced languages and complex tasks. In the EU21-GSM8K and EU21-MMLU tasks the median performance is the worst compared to the other models.
Qwen2-7B-Instruct demonstrates the highest median accuracy in EU21-MMLU, but performs worse in EU21-Hellaswag and EU21-ARC. In EU21-GSM8k the model performs similar to Mistral-NeMo-Minitron-8B-Base and Meta-Llama-3.1-8B-Instruct.
Medium-Sized Models ($>$9B & $<$27B Parameters)
Figure 2: Accuracy results for four medium sized LLMs on the EU21 tasks.
This analysis (cf. Figure 2) focuses on medium-sized LLMs such as Mistral-Nemo-Instruct-12.2B [^12] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024), Phi-3-medium-14B-4k-Instruct Abdin et al. (2024) and Vicuna-13b-v1.5 Peng et al. (2023).
Mixtral-8x7B-Instruct-v0.1 shows the highest median performance on four out of the five tasks. Only on EU21-GSM8K its performance is lacking behind Mistral-Nemo-Instruct-12.2B, indicating worse performance on math related tasks. However, its performance tends to be more variable than Mistral-Nemo-Instruct-12.2B_2407, with some languages like Latvian displaying significantly lower scores, reflecting higher inconsistency across the board.
Mistral-Nemo-Instruct-12.2B_2407 delivers solid results across all tasks and languages, with a particularly strong median performance in the EU21-GSM8K task. As evidenced by its narrow interquartile ranges the model signals low variability across all 21 languages compared to the other models.
Phi-3-medium-14B-4k-Instruct presents competitive results mostly in English by having the highest GSM8K, MMLU values but lags vastly behind the previous two models in median peformance across all languages and tasks, indicating only a small amount of multilingual training corpora. Lastly, Vicuna-13-b-v1.5 based on the older LLama-2-13B is not competitive on any task, due to its missing multilingual capabilities.
Compared to the small-sized category, Mixtral-8x7B-Instruct-v0.1 provides a substantial gain in EU21-Hellswag by achieving an median performance of 0.64, an absolute increase of 4% compared to Gemma-2-9b-Instruct. On EU21-GSM8K Mistral-Nemo-Instruct-12.2B achieves 57% accuracy a small increase of 1% compared to Meta-Llama-3.1-8B-Instruct. On EU21-MMLU Qwen2-7B achieved a accuracy of 59.4%, slightly behind medium sized models such as Mistral-Nemo-Base-12.2B with 60.2% and Mixtral-8x7B-Instruct-v0.1 with an performance of 60.8%.
Overall the medium category provides only small task dependend gains with specific models compared to the smaller Gemma-2-9b-Instruct model, indicating its strength in multilinguality.
Large-Sized Models ($\geq$27B Parameters)
Figure 3: Accuracy results for four large sized LLMs on the EU21 tasks.
This analysis (cf. Figure 3) focuses on large-sized LLMs such as Gemma-2-27b-Instruct Team et al. (2024b), Vicuna-33b-v1.3 Peng et al. (2023), Meta-Llama-3.1-70B-Instruct Dubey et al. (2024) and c4ai-command-r-35B-v01131313https://huggingface.co/CohereForAI/c4ai-command-r-v01. Gemma-2-27b-Instruct and Meta-Llama-3.1-70B-Instruct reliable outperforms the other models across most tasks, with median scores often exceeding an accuracy of 0.7. Their performance is not only high but also stable across different languages, as evidenced by its narrow interquartile ranges, signaling low variability. Probably due to the higher capacity (compared to Gemma-2-9b), Gemma-2-27b-Instruct can now also score higher on specialized knowledge tasks such as MMLU and GSM8K.
Vicuna-33b-v1.3 and c4ai-command-r-35B-v01 perform worse compared to the previous two models. This indicates a monolingual training corpora paired with a monolingual tokenizer.
Especially Gemma-2-27b-Instruct convinces with a remarkable performance. It provides a vast improvements over the small and medium sized LLMs and is even competitive with the more than twice as large Meta-LLama-3.1 70B model.
4.2 Impact of available training corpora
To better analyze the performance of LLMs across different languages, we categorized our target languages into high-resource (HRL) and medium-resource (MRL) groups, based on their proportion in the CommonCrawl dataset, which serves as the primary pre-training source for many LLMs. High-resource languages (HRL) had a CommonCrawl representation of 1% or more, whereas medium-resource languages (MRL) had between 0.01% and 1%. The language statistics accompanied with evaluation results of the LLama-3.1-70B-Instruct Dubey et al. (2024) model is depicted in Table 1.
As can be seen there is a downwards trend with outliers in Swedish and Danish, both Germanic languages that might profit from English or other Germanic training corpora (cf. Table 1).
Lang. Speak. M [^14] https://www.ethnologue.com/insights/ethnologue200/ CC(%) [^15] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages Cat. Avg. ARC GSM8K HeSw MMLU TQA
English 1456 45.05 HRL .79 .80 .87 .86 .82 .60
Spanish 559 4.50 HRL .73 .76 .73 .79 .80 .56
French 310 4.22 HRL .72 .75 .75 .78 .78 .56
Port. 264 2.02 HRL .73 .73 .78 .78 .78 .57
German 133 5.47 HRL .71 .73 .74 .75 .78 .57
Italian 68 2.64 HRL .73 .76 .76 .77 .79 .58
Polish 40.6 1.68 HRL .70 .72 .69 .74 .78 .59
Dutch 24.5 1.83 HRL .73 .73 .80 .76 .78 .60
Czech 12.3 1.09 HRL .71 .72 .75 .72 .78 .58
Romanian 24.5 0.52 MRL .72 .73 .73 .74 .79 .59
Greek 13.1 0.55 MRL .70 .69 .78 .71 .74 .58
Swedish 13.1 0.63 MRL .75 .74 .81 .78 .78 .62
Hungarian 12.6 0.56 MRL .70 .70 .75 .72 .77 .59
Bulgarian 7.7 0.29 MRL .66 .76 .76 .70 .67 .62
Slovak 7.3 0.37 MRL .69 .75 .74 .70 .68 .60
Danish 5.6 0.45 MRL .73 .72 .79 .75 .79 .59
Finnish 5.6 0.35 MRL .68 .67 .77 .71 .75 .52
Lithuanian 3.8 0.16 MRL .65 .72 .72 .66 .64 .60
Slovenian 2.9 0.15 MRL .67 .68 .72 .66 .74 .55
Latvian 1.5 0.12 MRL .66 .64 .71 .61 .73 .58
Estonian 1.1 0.14 MRL .66 .64 .76 .64 .74 .52
Table 1: Language statistics table sorted by number of speakers (approximate total speakers worldwide), and the accuracy score for the LLama-3.1 70B-Instruct model in each task and language.
4.3 Influence of Language Family
In this subsection we investigate how the LLM performance depend on the language family. We categorize the larger language groups based on their corresponding language family like the following:
  • Germanic: Danish, German, English, Dutch, Swedish (in total 53.38% of CC)
  • Romance: Spanish, French , Italian, Portuguese, Romanian (in total 13.9% of CC)
  • Slavic: Bulgarian, Czech, Polish, Slovak, Slovenian (in total 3.58% of CC)
  • Not categorized: Greek (hellenic), Hungarian (ugric), Estonian (finnic), Finnish (finnic), Lithuanian (baltic), Latvian (baltic)
Examining the average accuracies (cf. Appendix LABEL:tab:acc_std_lang_group_germanic_1, LABEL:tab:acc_std_lang_group_germanic_2, LABEL:tab:acc_std_lang_group_romance_1, LABEL:tab:acc_std_lang_group_romance_2 LABEL:tab:acc_std_lang_group_slavic_1, LABEL:tab:acc_std_lang_group_slavic_2) across the three language families, we observe that models generally achieve higher performance on Romance and Germanic languages compared to Slavic languages. For instance, the Gemma-2-27b-InstructTeam et al. (2024b) model attains average accuracies of 0.727 on Germanic languages (cf. LABEL:tab:acc_std_lang_group_germanic_1), 0.716 on Romance languages (cf. LABEL:tab:acc_std_lang_group_romance_1), and 0.694 on Slavic languages (cf. LABEL:tab:acc_std_lang_group_slavic_1).
Similarly, the Meta-Llama-3.1-70B-Instruct Dubey et al. (2024) model achieves average accuracies of 0.745 (Germanic), 0.726 (Romance), and 0.680 (Slavic), reinforcing the trend of lower performance on Slavic languages. In the following we compare the model performance on similar sized (in CC %) languages such as Romanian (Romance, 0.52%), Swedish (Germanic, 0.63%) and Polish (Slavic, 1.09%). Table 1 indicates that even though, Swedish is less represented than Polish, it achieves higher performance across all tasks. Also Romanian as the smallest represented language achieves higher performance across all five tasks, indicating that the language family is of influence in the performance analysis.
5 Correlation with Human Preference Scores
We examine the correlation between our introduced EU21 benchmarks and the Elo scores from the LMSYS Chatbot Arena [^16] https://chat.lmsys.org/?leaderboard to better understand how human preferences are reflected.
The Elo scores we use are derived from a dataset [^17] https://storage.googleapis.com/arena_external_data/public/clean_battle_20240814_public.json of about 1.8 million human quality judgements. In the LMSYS Chatbot Arena Chiang et al. (2024) users can vote on a pair of LLM responses in anonymized and randomized "battles".
The battles are divided into several categories, some of which correspond to the response languages. The Elo rankings are computed independently for each category.
In order to quantify how well the accuracies of a given model on our translated datasets correlate with the Elo scores, we compute the Pearson and Spearman Correlation Coefficients between the accuracy and the Elo scores of the $n=17$ different models. We compute these correlation separately for the Czech, French, German, Italian, Polish, Portuegese, Spanish and for English as a reference point. The remaining mostly lower ressource languages have a high uncertainty on their Elo scores as fewer battles (cf. Table 20) are available.
The LLMs for the analysis were mostly chosen based on the Elo rankings available, according to the following criteria:
  • Text-only, Causal Transformer-based LM
  • Checkpoints available on HF Hub
  • Model Size below 100B parameters
  • LMSYS Elo Ratings available for English, French, German and Spanish
Lang. Pearson Spearman
CS .570** .430*
EN .716*** .618**
FR .770*** .745***
DE .775*** .740***
IT .694*** .674***
PL .895*** .858***
PT .811*** .806***
ES .751*** .784***
Table 2: Correlation Coefficients for Task Avg. between Few-Shot Accuracy and LMSYS Elo Score. Significance levels: *p < 0.05, **p < 0.01, ***p < 0.001.
[…]
6 Influence of the translation service
While OKAPI Lai et al. (2023) and our evaluation datasets share a similar structure, they differ in key aspects such as translation methodology and sample alignment. The most notable difference between the two datasets is the translation approach: The OGX dataset was translated using DeepL, whereas OKAPI employed ChatGPT for this task. The translations for the OKAPI tasks cover 31 languages, including both EU and non-EU languages. For the OpenGPT-X evaluation datasets, we focused on translations into 20 European languages. The common and new set of EU languages between the OKAPI and OpenGPT-X are shown in Table 3.
Common Languages New Languages
Danish Czech
German Greek
Spanish Estonian
French Finnish
Hungarian Lithuanian
Italian Latvian
Dutch Bulgarian
Romanian Polish
Slovak Slovenian
Swedish
Table 3: Comparison of languages between OKAPI and OpenGPT-X Datasets
In this section, we present a correlation analysis with a focus on the ARC, HellaSwag, and MMLU tasks, alongside a COMET score Rei et al. (2023) evaluation to assess the translation quality across both datasets.
The idea behind the correlation analysis is to record [^18] The evaluation was performed using the LM-Evaluation-Harness framework, with sample-level logging enabled.whether the model responded correctly or incorrectly for each OKAPI and OGX sample and to examine if the predictions are correlated, allowing direct performance comparison on sample-level on the same data. For a sample-level comparison, we ensured that OGX and OKAPI samples are matched and can be linked to their original translations. While ARC, HellaSwag, and MMLU samples have IDs matching the original English subsets, enabling mapping, this does not apply to TruthfulQA.
For the correlation analysis, we calculated the correlation matrices for different models, tasks, and languages, listed in Appendix A.
The highest correlations, as shown in Table 17, are with Aya-23-8B (0.57), Meta-Llama-3-8B-Instruct (0.56), and gemma-7b (0.56), while bloom-7b1 (0.40) and bloomz-7b1 (0.44) show lower correlations. This indicates that Aya and Meta-Llama perform more consistently, while bloom-7b1 and bloomz-7b1 exhibit greater variability.
As seen in Table 19, ARC has the highest correlation at 0.61, followed by Hellaswag and MMLU (both 0.54). ARC’s structure and reliance on factual knowledge likely make model performance more stable and predictable.
Regarding the task performance correlation, we can see from Table 18 that correlations range from 0.44 for Hungarian to 0.59 for Spanish. Hungarian’s lower score may reflect fewer training resources, while Spanish likely benefits from better data availability, leading to more consistent model performance.
Translation quality likely plays a role in these results, as ambiguous or incorrect translations, poorly adapted prompts, or task mismatches can negatively affect model performance. Thus, the observed correlations may also reflect the degree to which translation quality impacts each language’s evaluation results. Table 5, Table 6, Fig. 4 and Table 7 The examples in Appendix A Table 5, Table 6, Fig. 4 and Table 7 underscore the importance of context-aware translation in multilingual tasks, as naive translations can reduce clarity and coherence, ultimately affecting model performance.
As part of the further analysis, we determined the contingency table (4) to compare the answers of the models on our evaluation datasets and the OKAPI datasets. Both evaluations show similar alignment, with agreements on correct answers (38.1%) and incorrect answers (39.3%) being fairly balanced. However, OKAPI seems to perform slightly better in cases of disagreement, as OKAPI is correct and our model is incorrect in 11.8% of cases, compared to 10.8% where our model is correct and OKAPI is incorrect. This suggests that, while the performance of both models is comparable, OKAPI might have a slight edge in terms of accuracy in cases where they disagree.
OKAPI Corr. OKAPI Inc.
Ours Corr. 0.3808 0.1183
Ours Inc. 0.1077 0.3929
Table 4: Contingency table comparing the answers of the models on the evaluations performed on our translations and the OKAPI translations.
Further we used the COMET score Rei et al. (2023) to analyze the translation quality of both datasets. Since there is no target translation for the different benchmarks, we evaluated the translation quality of the OGX and OKAPI datasets using the COMET-KIWI metric Rei et al. (2023). This metric uses a pre-trained language model to predict translation quality. A score of 1 indicates a high-quality translation and a score close to 0 indicates a bad translation. Both the OGX and OKAPI datasets have a relatively high average score of 0.8208 and 0.8190, respectively, across all tasks and translations. The exact scores can be found in the Appendix in Table Table 29. It is important to highlight that because there is no target translation, no statistical significance can be calculated between the scores of OKAPI and OGX. Therefore, we cannot conclude that the OGX translations are automatically higher quality based on the COMET KIWI score. Nevertheless, it can be shown that the OGX datasets have a consistently high translation quality across all languages (See Appendix in Table Table 29).
We additionally compared all our models on the ARC benchmark Fig. 10, Hellaswag Fig. 11 and MMLU Fig. 12. Overall we can see that all evaluated models achieve higher accuracies on our tasks compared to OKAPI. This analysis highlights the extent to which our and OKAPI results align across models, tasks, and languages, providing insights into the consistency and reliability of model evaluations across different datasets.
7 Conclusion
This work provides a comprehensive framework for evaluating Large Language Models (LLMs) across multiple European languages by utilizing translated versions of widely-used benchmarks. Our experiments demonstrate that translated benchmarks can serve as a reliable proxy for human preference in various languages. We also highlight the influence of translation services on the quality of the benchmarks. By releasing our benchmarks, we aim to foster further research and development in multilingual LLM evaluation, driving improvements in cross-lingual NLP applications. Future work will focus on refining cultural localization in translations and exploring deeper correlations between translation quality and model performance across diverse language families.
8 Limitations
A hindrance in achieving benchmarks comparable across languages is localization of cultural aspects. Toxicity benchmarks in particular can suffer from this, as stereotypes and terms related to them likely do not translate directly been cultures and hence languages. To a lesser extent this applies to all benchmarks requiring some degree of common sense and world knowledge. A team of reseachers has launched an initiative [^19] https://huggingface.co/spaces/CohereForAI/MMLU-evaluationto correct some of these cultural biases in multilingual datasets (predominantly MMLU), through crowd-sourced manual work. While we are aware of this initiative, we opted not to conduct a similar enterprise in order to avoid doubled efforts. Aligning the results from the initative with our own translations will be an avenue for further work.
Furthermore the degree to which a translation for an given task is not always feasible. The pronoun-resolution benchmark WinoGrande Sakaguchi et al. (2019) for instance is posed as a multiple choice task, where a on object or subject in a sentence is masked and each choice constitutes a replacement of the mask. The model is then tasked with selecting the sentence that is semantically correct. Tables LABEL:fig:winogrande_base and LABEL:fig:winogrande_spa demonstrate how in languages with gendered nouns or gendered adjectives in predicative use, the benchmark becomes easier due to strong grammatical signals indicating the correct choice of subject or object. It is for that reason we decided against translation of WinoGrande.
In order to compare the translations of OGX with the original benchmark or other translations such as OKAPI, the samples must be mapped. This requires a one-to-one mapping. However, this mapping was impossible for all benchmarks; therefore, we could only compare three benchmarks with the OKAPI dataset: Hellaswag, MMLU, and ARC. For Hellaswag, we used the ind and the source_id; for ARC, we used the id; and for MMLU, the subtask_name and the natural sorting of the benchmarks, which turned out to be a one-to-one mapping.
We analyzed the translated sequences with unexpectedly low scores (<0.5) during the quality analysis of the translations using the COMET KIWI score. Depending on the benchmark, some sequences were not translated. As a result, some answer options are either in the original language or do not appear at all. The maximum error rate is 2.3 % for Hellaswag, 1.8 % for ARC, and 2.1 % for MMLU across all languages. Although the errors show no significant difference in direct comparison with OKAPI (see Section 6) nor with the original benchmarks, the analysis using the COMET score in the translation pipeline should be considered in future work.
No measures to indicate overfitting on English benchmarks, such as math-based benchmarks as in https://arxiv.org/pdf/2405.00332
Appendix A Experimental Setup
Context: A cartoon animation video is shown with people wandering around and rockets being shot. two men
Endings:
1. fight robots of evil and ends with a to be continued.
2. are then shown in closeups shooting a shot put.
3. push a child in a speedboat in the water.
4. look in the cameraman’s eye and smile.
Table 5: An English HellaSwag sample, which includes a context (a cartoon animation video) and four possible endings. The task requires models to select the most appropriate ending based on the context.
Context: Es wird ein Zeichentrickvideo gezeigt, in dem Menschen umherwandern und Raketen geschossen werden. zwei Männer
Endings:
1. Kampfroboter des Bösen und endet mit einem "to be continued"
2. werden dann in Nahaufnahmen beim Kugelstoßen gezeigt.
3. ein Kind in einem Schnellboot ins Wasser schieben.
4. in die Augen des Kameramanns schauen und lächeln.
Table 6: Naive, context-unaware translation into German of the sample shown in Table 5. This literal translation demonstrates that without considering the context, certain phrases such as "Raketen geschossen" (rockets shot) lose fluency and clarity in translation, affecting the overall coherence of the text.
    ’A cartoon animation video is shown
    with people wandering around and
    rockets being shot. two men
    <x>SEP</x>fight robots
    of evil and ends with a to be
    continued.<x>SEP</x>are then shown
    in closeupsshooting a shot put.
    <x>SEP</x>look in the cameraman’s
    eye and smile.’
Figure 4: Processed HellaSwag instance, as supplied to the DeepL-API.The original context and endings are formatted in a machine-readable way, ensuring consistent input for translation.
Context: Es wird ein Zeichentrickfilm gezeigt, in dem Menschen umherlaufen und Raketen abgeschossen werden. zwei Männer’
Endings:
1. kämpfen gegen Roboter des Bösen und enden mit einem "to be continued".
2. werden dann in Großaufnahme beim Kugelstoßen gezeigt.
3. schieben ein Kind in einem Schnellboot im Wasser.
4. schauen dem Kameramann in die Augen und lächeln.
Table 7: Context-aware translation into German of the sample shown in Table 5. In contrast to the naive translation, this version properly adapts phrases, such as translating "rockets being shot" to "Raketen abgeschossen" (rockets launched), maintaining fluency and better capturing the meaning of the original text.
Appendix B European LLM Evaluation Results
Table 8: Performance of Models Across Tasks: Accuracy and Standard Deviation for 21 Languages
Task Group
Model Name Avg. ARC HellaSwag
Aya-23-8B Aryabumi et al. (2024) .441 .475 ± .134 .535 ± .145
Bloom-7b1 Workshop et al. (2023) .280 .319 ± .098 .355 ± .112
Bloomz-7b1 Muennighoff et al. (2023) .287 .316 ± .101 .354 ± .115
EuroLLM-1.7B Martins et al. (2024) .338 .469 ± .036 .485 ± .038
EuroLLM-1.7B-Instruct Martins et al. (2024) .358 .486 ± .038 .495 ± .039
Gemma-1.1-7b-Instruct Team et al. (2024a) .373 .443 ± .072 .437 ± .085
Gemma-2-27b-Instruct Yang et al. (2024) .698 .747 ± .034 .711 ± .048
Gemma-2-9b-Instruct Yang et al. (2024) .581 .668 ± .061 .608 ± .069
Gemma-7b Team et al. (2024a) .536 .607 ± .052 .629 ± .066
Meta-Llama-2-13B-Chat Touvron et al. (2023) .407 .462 ± .116 .504 ± .125
Meta-Llama-2-7B Touvron et al. (2023) .362 .447 ± .108 .482 ± .114
Meta-Llama-2-7B-Chat Touvron et al. (2023) .370 .444 ± .110 .457 ± .114
Meta-Llama-3-8B Dubey et al. (2024) .503 .550 ± .074 .587 ± .089
Meta-Llama-3-8B-Instruct Dubey et al. (2024) .547 .558 ± .080 .539 ± .091
Meta-Llama-3.1-70B-Instruct Dubey et al. (2024) .706 .712 ± .040 .728 ± .059
Meta-Llama-3.1-8B Dubey et al. (2024) .512 .554 ± .071 .588 ± .090
Meta-Llama-3.1-8B-Instruct Dubey et al. (2024) .562 .563 ± .075 .579 ± .089
Mistral-7B-Instruct-v0.1 Jiang et al. (2023) .400 .446 ± .121 .454 ± .113
Mistral-7B-Instruct-v0.2 Jiang et al. (2023) .475 .524 ± .135 .539 ± .133
Mistral-7B-Instruct-v0.3 Jiang et al. (2023) .482 .530 ± .136 .538 ± .129
Mistral-7B-v0.1 Jiang et al. (2023) .458 .521 ± .130 .539 ± .125
Mistral-7B-v0.3 Jiang et al. (2023) .450 .513 ± .128 .534 ± .124
Mistral-NeMo-Minitron-8B-Base Sreenivas et al. (2024) .502 .522 ± .114 .546 ± .120
Mistral-Nemo-Base-12.2B_2407202020https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 .562 .615 ± .071 .641 ± .094
Mistral-Nemo-Instruct-12.2B_2407212121https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 .598 .621 ± .071 .624 ± .093
Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024) .591 .625 ± .138 .641 ± .137
Mixtral-8x7B-v0.1 Jiang et al. (2024) .554 .608 ± .132 .636 ± .132
Occiglot-7b-eu5222222https://huggingface.co/occiglot/occiglot-7b-eu5 .404 .470 ± .130 .511 ± .143
Occiglot-7b-eu5-Instruct [^23] https://huggingface.co/occiglot/occiglot-7b-eu5-instruct .406 .484 ± .132 .519 ± .146
Pharia-1-LLM-7B-control  [^24] https://huggingface.co/Aleph-Alpha/Pharia-1-LLM-7B-control .333 .393 ± .174 .433 ± .181
Pharia-1-LLM-7B-control-aligned [^25] https://huggingface.co/Aleph-Alpha/Pharia-1-LLM-7B-control-aligned .346 .396 ± .177 .438 ± .185
Phi-3-medium-14B-128k-Instruct Abdin et al. (2024) .508 .497 ± .168 .492 ± .166
Phi-3-medium-14B-4k-Instruct Abdin et al. (2024) .519 .500 ± .170 .500 ± .169
Phi-3-mini-3.8B-128k-Instruct Abdin et al. (2024) .424 .415 ± .161 .420 ± .153
Phi-3-mini-4k-Instruct Abdin et al. (2024) .415 .411 ± .161 .422 ± .158
Qwen2-7B Yang et al. (2024) .542 .481 ± .113 .519 ± .116
Qwen2-7B-Instruct Yang et al. (2024) .528 .506 ± .115 .530 ± .120
Vicuna-13b-v1.5 Peng et al. (2023) .426 .467 ± .111 .519 ± .124
Vicuna-33b-v1.3 Peng et al. (2023) .447 .527 ± .140 .547 ± .135
c4ai-command-r-35B-v01262626https://huggingface.co/CohereForAI/c4ai-command-r-v01 .557 .592 ± .130 .646 ± .135
Table 9: Performance of Models Across Tasks: Accuracy and Standard Deviation for 21 Languages
Task Group
Model Name MMLU TruthfulQA GSM8k
Aya-23-8B .455 ± .070 .476 ± .028 .266 ± .105
Bloom-7b1 .256 ± .005 .464 ± .039 .010 ± .005
Bloomz-7b1 .302 ± .039 .461 ± .037 .000 ± .001
EuroLLM-1.7B .257 ± .006 .441 ± .027 .035 ± .006
EuroLLM-1.7B-Instruct .266 ± .007 .462 ± .028 .082 ± .021
Gemma-1.1-7b-Instruct .414 ± .044 .466 ± .026 .106 ± .049
Gemma-2-27b-Instruct .679 ± .027 .605 ± .021 .748 ± .023
Gemma-2-9b-Instruct .590 ± .042 .588 ± .022 .448 ± .042
Gemma-7b .568 ± .032 .479 ± .027 .394 ± .036
Meta-Llama-2-13B-Chat .429 ± .063 .469 ± .028 .169 ± .081
Meta-Llama-2-7B .364 ± .050 .438 ± .031 .079 ± .033
Meta-Llama-2-7B-Chat .370 ± .052 .478 ± .029 .100 ± .052
Meta-Llama-3-8B .554 ± .049 .477 ± .027 .346 ± .058
Meta-Llama-3-8B-Instruct .548 ± .055 .532 ± .024 .560 ± .085
Meta-Llama-3.1-70B-Instruct .771 ± .026 .573 ± .026 .746 ± .063
Meta-Llama-3.1-8B .556 ± .047 .495 ± .026 .365 ± .057
Meta-Llama-3.1-8B-Instruct .576 ± .051 .532 ± .023 .560 ± .071
Mistral-7B-Instruct-v0.1 .424 ± .070 .512 ± .051 .165 ± .095
Mistral-7B-Instruct-v0.2 .476 ± .072 .585 ± .066 .251 ± .089
Mistral-7B-Instruct-v0.3 .491 ± .075 .548 ± .048 .303 ± .101
Mistral-7B-v0.1 .506 ± .077 .475 ± .035 .248 ± .082
Mistral-7B-v0.3 .501 ± .075 .472 ± .034 .231 ± .076
Mistral-NeMo-Minitron-8B-Base .555 ± .069 .481 ± .026 .407 ± .076
Mistral-Nemo-Base-12.2B_2407 .602 ± .047 .513 ± .031 .441 ± .058
Mistral-Nemo-Instruct-12.2B_2407 .595 ± .047 .577 ± .027 .571 ± .071
Mixtral-8x7B-Instruct-v0.1 .608 ± .080 .605 ± .048 .477 ± .108
Mixtral-8x7B-v0.1 .615 ± .079 .488 ± .030 .424 ± .102
Occiglot-7b-eu5 .426 ± .073 .448 ± .030 .163 ± .065
Occiglot-7b-eu5-Instruct .428 ± .074 .471 ± .032 .131 ± .055
Pharia-1-LLM-7B-control .353 ± .080 .456 ± .029 .030 ± .029
Pharia-1-LLM-7B-control-aligned .366 ± .088 .469 ± .034 .060 ± .065
Phi-3-medium-14B-128k-Instruct .565 ± .123 .494 ± .047 .491 ± .206
Phi-3-medium-14B-4k-Instruct .574 ± .127 .499 ± .053 .523 ± .202
Phi-3-mini-3.8B-128k-Instruct .462 ± .125 .501 ± .043 .320 ± .244
Phi-3-mini-4k-Instruct .449 ± .118 .498 ± .052 .292 ± .230
Qwen2-7B .593 ± .066 .527 ± .025 .592 ± .105
Qwen2-7B-Instruct .584 ± .065 .553 ± .030 .468 ± .116
Vicuna-13b-v1.5 .448 ± .067 .505 ± .030 .188 ± .071
Vicuna-33b-v1.3 .467 ± .078 .524 ± .039 .169 ± .082
c4ai-command-r-35B-v01 .561 ± .074 .543 ± .032 .440 ± .085
Table 10: Standard Deviation of Task Accuracies for Germanic Languages
Task Group
Model Name Avg. ARC HellaSwag
Aya-23-8B .509 .563 ± .105 .636 ± .121
Bloom-7b1 .291 .348 ± .132 .381 ± .150
Bloomz-7b1 .303 .344 ± .138 .379 ± .154
EuroLLM-1.7B .352 .507 ± .046 .518 ± .058
EuroLLM-1.7B-Instruct .377 .520 ± .051 .530 ± .058
Gemma-1.1-7b-Instruct .422 .509 ± .090 .517 ± .134
Gemma-2-27b-Instruct .727 .776 ± .037 .757 ± .060
Gemma-2-9b-Instruct .624 .723 ± .053 .674 ± .082
Gemma-7b .574 .655 ± .053 .694 ± .084
Meta-Llama-2-13B-Chat .485 .566 ± .079 .635 ± .125
Meta-Llama-2-7B .420 .544 ± .087 .599 ± .125
Meta-Llama-2-7B-Chat .438 .542 ± .089 .578 ± .140
Meta-Llama-3-8B .559 .615 ± .069 .679 ± .095
Meta-Llama-3-8B-Instruct .614 .627 ± .076 .636 ± .102
Meta-Llama-3.1-70B-Instruct .745 .748 ± .037 .787 ± .052
Meta-Llama-3.1-8B .568 .618 ± .064 .682 ± .092
Meta-Llama-3.1-8B-Instruct .624 .633 ± .069 .672 ± .087
Mistral-7B-Instruct-v0.1 .483 .552 ± .088 .566 ± .128
Mistral-7B-Instruct-v0.2 .569 .634 ± .072 .667 ± .120
Mistral-7B-Instruct-v0.3 .571 .642 ± .075 .663 ± .122
Mistral-7B-v0.1 .534 .629 ± .072 .658 ± .119
Mistral-7B-v0.3 .522 .616 ± .075 .652 ± .118
Mistral-NeMo-Minitron-8B-Base .577 .620 ± .093 .657 ± .119
Mistral-Nemo-Base-12.2B_2407 .614 .671 ± .068 .724 ± .091
Mistral-Nemo-Instruct-12.2B_2407 .654 .680 ± .068 .708 ± .094
Mixtral-8x7B-Instruct-v0.1 .676 .727 ± .054 .760 ± .080
Mixtral-8x7B-v0.1 .632 .706 ± .050 .751 ± .079
Occiglot-7b-eu5 .475 .577 ± .087 .637 ± .121
Occiglot-7b-eu5-Instruct .473 .591 ± .087 .650 ± .117
Pharia-1-LLM-7B-control .428 .558 ± .165 .605 ± .181
Pharia-1-LLM-7B-control-aligned .450 .562 ± .167 .612 ± .182
Phi-3-medium-14B-128k-Instruct .633 .634 ± .129 .634 ± .166
Phi-3-medium-14B-4k-Instruct .646 .637 ± .129 .644 ± .167
Phi-3-mini-3.8B-128k-Instruct .545 .540 ± .174 .538 ± .193
Phi-3-mini-4k-Instruct .536 .543 ± .170 .545 ± .198
Qwen2-7B .622 .580 ± .111 .629 ± .121
Qwen2-7B-Instruct .597 .607 ± .101 .644 ± .118
Vicuna-13b-v1.5 .502 .566 ± .066 .645 ± .111
Vicuna-33b-v1.3 .532 .644 ± .072 .678 ± .103
c4ai-command-r-35B-v01 .627 .681 ± .076 .750 ± .087
Table 11: Standard Deviation of Task Accuracies for Germanic Languages
Task Group
Model Name MMLU TruthfulQA GSM8k
Aya-23-8B .509 ± .050 .490 ± .026 .343 ± .077
Bloom-7b1 .257 ± .004 .450 ± .051 .016 ± .005
Bloomz-7b1 .321 ± .034 .471 ± .028 .001 ± .001
EuroLLM-1.7B .261 ± .007 .432 ± .036 .040 ± .005
EuroLLM-1.7B-Instruct .272 ± .007 .458 ± .036 .106 ± .030
Gemma-1.1-7b-Instruct .453 ± .054 .467 ± .016 .163 ± .075
Gemma-2-27b-Instruct .704 ± .024 .618 ± .020 .778 ± .004
Gemma-2-9b-Instruct .630 ± .038 .601 ± .021 .492 ± .043
Gemma-7b .599 ± .029 .486 ± .022 .438 ± .041
Meta-Llama-2-13B-Chat .490 ± .038 .481 ± .030 .251 ± .058
Meta-Llama-2-7B .411 ± .039 .436 ± .035 .108 ± .023
Meta-Llama-2-7B-Chat .422 ± .041 .497 ± .033 .150 ± .043
Meta-Llama-3-8B .603 ± .042 .486 ± .040 .414 ± .059
Meta-Llama-3-8B-Instruct .601 ± .047 .551 ± .025 .656 ± .070
Meta-Llama-3.1-70B-Instruct .789 ± .019 .599 ± .023 .804 ± .055
Meta-Llama-3.1-8B .604 ± .038 .507 ± .044 .431 ± .047
Meta-Llama-3.1-8B-Instruct .627 ± .042 .550 ± .025 .636 ± .079
Mistral-7B-Instruct-v0.1 .488 ± .046 .555 ± .013 .254 ± .051
Mistral-7B-Instruct-v0.2 .542 ± .042 .657 ± .012 .343 ± .050
Mistral-7B-Instruct-v0.3 .557 ± .047 .595 ± .019 .399 ± .040
Mistral-7B-v0.1 .575 ± .041 .486 ± .048 .323 ± .043
Mistral-7B-v0.3 .568 ± .044 .478 ± .040 .295 ± .039
Mistral-NeMo-Minitron-8B-Base .622 ± .045 .497 ± .023 .487 ± .067
Mistral-Nemo-Base-12.2B_2407 .646 ± .036 .529 ± .038 .500 ± .045
Mistral-Nemo-Instruct-12.2B_2407 .639 ± .037 .587 ± .041 .654 ± .060
Mixtral-8x7B-Instruct-v0.1 .670 ± .029 .646 ± .002 .576 ± .040
Mixtral-8x7B-v0.1 .678 ± .028 .507 ± .019 .519 ± .043
Occiglot-7b-eu5 .491 ± .044 .449 ± .035 .222 ± .037
Occiglot-7b-eu5-Instruct .496 ± .045 .483 ± .033 .144 ± .020
Pharia-1-LLM-7B-control .435 ± .069 .477 ± .025 .067 ± .034
Pharia-1-LLM-7B-control-aligned .458 ± .069 .492 ± .015 .128 ± .063
Phi-3-medium-14B-128k-Instruct .675 ± .078 .533 ± .018 .687 ± .120
Phi-3-medium-14B-4k-Instruct .688 ± .077 .549 ± .026 .715 ± .109
Phi-3-mini-3.8B-128k-Instruct .575 ± .102 .532 ± .019 .540 ± .205
Phi-3-mini-4k-Instruct .559 ± .103 .534 ± .048 .500 ± .216
Qwen2-7B .654 ± .045 .547 ± .012 .698 ± .057
Qwen2-7B-Instruct .646 ± .047 .585 ± .014 .504 ± .190
Vicuna-13b-v1.5 .508 ± .038 .533 ± .022 .257 ± .026
Vicuna-33b-v1.3 .537 ± .038 .570 ± .025 .233 ± .064
c4ai-command-r-35B-v01 .620 ± .044 .570 ± .017 .512 ± .053
Table 12: Standard Deviation of Task Accuracies for Romance Languages
Task Group
Model Name Avg. ARC HellaSwag
Aya-23-8B .522 .600 ± .013 .668 ± .019
Bloom-7b1 .323 .427 ± .103 .478 ± .125
Bloomz-7b1 .337 .427 ± .106 .480 ± .130
EuroLLM-1.7B .346 .495 ± .019 .511 ± .024
EuroLLM-1.7B-Instruct .368 .515 ± .020 .524 ± .025
Gemma-1.1-7b-Instruct .403 .495 ± .031 .488 ± .038
Gemma-2-27b-Instruct .716 .772 ± .011 .743 ± .015
Gemma-2-9b-Instruct .617 .719 ± .018 .658 ± .028
Gemma-7b .557 .648 ± .020 .672 ± .029
Meta-Llama-2-13B-Chat .461 .543 ± .031 .584 ± .045
Meta-Llama-2-7B .400 .527 ± .028 .557 ± .042
Meta-Llama-2-7B-Chat .414 .529 ± .033 .528 ± .042
Meta-Llama-3-8B .538 .611 ± .026 .649 ± .034
Meta-Llama-3-8B-Instruct .593 .628 ± .026 .601 ± .038
Meta-Llama-3.1-70B-Instruct .726 .745 ± .013 .772 ± .017
Meta-Llama-3.1-8B .548 .615 ± .022 .654 ± .033
Meta-Llama-3.1-8B-Instruct .604 .626 ± .021 .646 ± .033
Mistral-7B-Instruct-v0.1 .464 .542 ± .038 .528 ± .052
Mistral-7B-Instruct-v0.2 .540 .625 ± .030 .629 ± .048
Mistral-7B-Instruct-v0.3 .544 .627 ± .028 .623 ± .044
Mistral-7B-v0.1 .514 .614 ± .026 .624 ± .042
Mistral-7B-v0.3 .508 .606 ± .029 .619 ± .041
Mistral-NeMo-Minitron-8B-Base .568 .631 ± .057 .649 ± .074
Mistral-Nemo-Base-12.2B_2407 .608 .680 ± .043 .726 ± .061
Mistral-Nemo-Instruct-12.2B_2407 .644 .687 ± .035 .706 ± .059
Mixtral-8x7B-Instruct-v0.1 .662 .723 ± .028 .748 ± .045
Mixtral-8x7B-v0.1 .616 .702 ± .031 .736 ± .042
Occiglot-7b-eu5 .477 .596 ± .062 .646 ± .097
Occiglot-7b-eu5-Instruct .471 .614 ± .060 .656 ± .098
Pharia-1-LLM-7B-control .421 .561 ± .139 .603 ± .152
Pharia-1-LLM-7B-control-aligned .450 .570 ± .139 .614 ± .156
Phi-3-medium-14B-128k-Instruct .648 .677 ± .106 .655 ± .129
Phi-3-medium-14B-4k-Instruct .661 .683 ± .111 .665 ± .133
Phi-3-mini-3.8B-128k-Instruct .571 .584 ± .120 .568 ± .127
Phi-3-mini-4k-Instruct .558 .578 ± .118 .574 ± .131
Qwen2-7B .612 .584 ± .057 .617 ± .072
Qwen2-7B-Instruct .592 .615 ± .053 .632 ± .077
Vicuna-13b-v1.5 .480 .547 ± .029 .603 ± .046
Vicuna-33b-v1.3 .512 .627 ± .025 .641 ± .034
c4ai-command-r-35B-v01 .634 .712 ± .022 .766 ± .023
Table 13: Standard Deviation of Task Accuracies for Romance Languages
Task Group
Model Name MMLU TruthfulQA GSM8k
Aya-23-8B .515 ± .007 .476 ± .029 .351 ± .026
Bloom-7b1 .260 ± .003 .440 ± .051 .012 ± .003
Bloomz-7b1 .349 ± .035 .432 ± .040 .000 ± .001
EuroLLM-1.7B .256 ± .008 .428 ± .030 .038 ± .006
EuroLLM-1.7B-Instruct .267 ± .006 .448 ± .029 .085 ± .012
Gemma-1.1-7b-Instruct .449 ± .019 .475 ± .023 .108 ± .022
Gemma-2-27b-Instruct .701 ± .002 .610 ± .010 .752 ± .021
Gemma-2-9b-Instruct .626 ± .007 .608 ± .010 .474 ± .040
Gemma-7b .594 ± .006 .465 ± .033 .406 ± .016
Meta-Llama-2-13B-Chat .474 ± .013 .474 ± .007 .233 ± .037
Meta-Llama-2-7B .400 ± .013 .416 ± .025 .101 ± .018
Meta-Llama-2-7B-Chat .407 ± .020 .472 ± .014 .135 ± .036
Meta-Llama-3-8B .594 ± .013 .459 ± .021 .376 ± .021
Meta-Llama-3-8B-Instruct .591 ± .019 .540 ± .007 .602 ± .045
Meta-Llama-3.1-70B-Instruct .791 ± .007 .570 ± .014 .751 ± .021
Meta-Llama-3.1-8B .594 ± .010 .483 ± .021 .396 ± .023
Meta-Llama-3.1-8B-Instruct .616 ± .013 .542 ± .019 .592 ± .036
Mistral-7B-Instruct-v0.1 .476 ± .020 .544 ± .010 .230 ± .047
Mistral-7B-Instruct-v0.2 .527 ± .012 .617 ± .018 .302 ± .034
Mistral-7B-Instruct-v0.3 .542 ± .013 .568 ± .017 .358 ± .029
Mistral-7B-v0.1 .565 ± .012 .465 ± .031 .303 ± .025
Mistral-7B-v0.3 .555 ± .011 .473 ± .038 .284 ± .029
Mistral-NeMo-Minitron-8B-Base .615 ± .026 .481 ± .014 .462 ± .040
Mistral-Nemo-Base-12.2B_2407 .645 ± .018 .504 ± .020 .485 ± .033
Mistral-Nemo-Instruct-12.2B_2407 .637 ± .019 .575 ± .024 .615 ± .046
Mixtral-8x7B-Instruct-v0.1 .671 ± .014 .626 ± .005 .541 ± .037
Mixtral-8x7B-v0.1 .675 ± .011 .481 ± .026 .487 ± .041
Occiglot-7b-eu5 .493 ± .032 .436 ± .032 .214 ± .042
Occiglot-7b-eu5-Instruct .495 ± .037 .468 ± .029 .122 ± .066
Pharia-1-LLM-7B-control .433 ± .047 .459 ± .026 .049 ± .022
Pharia-1-LLM-7B-control-aligned .452 ± .056 .499 ± .013 .117 ± .055
Phi-3-medium-14B-128k-Instruct .689 ± .060 .544 ± .023 .677 ± .112
Phi-3-medium-14B-4k-Instruct .702 ± .060 .551 ± .031 .705 ± .116
Phi-3-mini-3.8B-128k-Instruct .590 ± .079 .546 ± .007 .567 ± .177
Phi-3-mini-4k-Instruct .567 ± .074 .548 ± .028 .521 ± .164
Qwen2-7B .650 ± .021 .539 ± .020 .669 ± .051
Qwen2-7B-Instruct .637 ± .022 .578 ± .012 .500 ± .080
Vicuna-13b-v1.5 .499 ± .013 .515 ± .022 .235 ± .014
Vicuna-33b-v1.3 .521 ± .010 .531 ± .012 .240 ± .024
c4ai-command-r-35B-v01 .627 ± .009 .566 ± .026 .500 ± .023
Table 14: Standard Deviation of Task Accuracies for Slavic Languages
Task Group
Model Name Avg. ARC HellaSwag
Aya-23-8B .412 .427 ± .097 .477 ± .105
Bloom-7b1 .259 .263 ± .007 .295 ± .002
Bloomz-7b1 .262 .259 ± .013 .294 ± .004
EuroLLM-1.7B .332 .455 ± .013 .472 ± .010
EuroLLM-1.7B-Instruct .354 .474 ± .013 .481 ± .012
Gemma-1.1-7b-Instruct .365 .427 ± .017 .417 ± .024
Gemma-2-27b-Instruct .694 .742 ± .016 .695 ± .020
Gemma-2-9b-Instruct .567 .658 ± .027 .586 ± .032
Gemma-7b .530 .600 ± .015 .607 ± .025
Meta-Llama-2-13B-Chat .392 .450 ± .052 .473 ± .045
Meta-Llama-2-7B .351 .435 ± .043 .449 ± .036
Meta-Llama-2-7B-Chat .355 .424 ± .046 .419 ± .043
Meta-Llama-3-8B .482 .528 ± .029 .541 ± .034
Meta-Llama-3-8B-Instruct .522 .533 ± .029 .493 ± .035
Meta-Llama-3.1-70B-Instruct .680 .699 ± .019 .696 ± .032
Meta-Llama-3.1-8B .492 .535 ± .026 .544 ± .036
Meta-Llama-3.1-8B-Instruct .541 .543 ± .026 .533 ± .037
Mistral-7B-Instruct-v0.1 .396 .442 ± .052 .432 ± .043
Mistral-7B-Instruct-v0.2 .478 .541 ± .050 .530 ± .050
Mistral-7B-Instruct-v0.3 .486 .544 ± .055 .528 ± .048
Mistral-7B-v0.1 .461 .536 ± .052 .528 ± .045
Mistral-7B-v0.3 .455 .528 ± .053 .523 ± .048
Mistral-NeMo-Minitron-8B-Base .475 .488 ± .043 .496 ± .039
Mistral-Nemo-Base-12.2B_2407 .546 .597 ± .024 .601 ± .029
Mistral-Nemo-Instruct-12.2B_2407 .581 .605 ± .030 .585 ± .033
Mixtral-8x7B-Instruct-v0.1 .598 .644 ± .050 .629 ± .052
Mixtral-8x7B-v0.1 .560 .626 ± .052 .625 ± .048
Occiglot-7b-eu5 .380 .443 ± .038 .457 ± .029
Occiglot-7b-eu5-Instruct .396 .457 ± .040 .464 ± .032
Pharia-1-LLM-7B-control .263 .263 ± .009 .305 ± .008
Pharia-1-LLM-7B-control-aligned .265 .263 ± .014 .307 ± .007
Phi-3-medium-14B-128k-Instruct .449 .423 ± .035 .402 ± .028
Phi-3-medium-14B-4k-Instruct .459 .423 ± .033 .406 ± .027
Phi-3-mini-3.8B-128k-Instruct .342 .318 ± .026 .329 ± .018
Phi-3-mini-4k-Instruct .330 .311 ± .030 .323 ± .015
Qwen2-7B .526 .451 ± .049 .481 ± .052
Qwen2-7B-Instruct .519 .474 ± .053 .489 ± .056
Vicuna-13b-v1.5 .414 .454 ± .054 .489 ± .044
Vicuna-33b-v1.3 .443 .544 ± .058 .531 ± .050
c4ai-command-r-35B-v01 .535 .566 ± .079 .603 ± .087
Table 15: Standard Deviation of Task Accuracies for Slavic Languages
Task Group
Model Name MMLU TruthfulQA GSM8k
Aya-23-8B .436 ± .047 .477 ± .030 .244 ± .056
Bloom-7b1 .255 ± .006 .475 ± .030 .007 ± .001
Bloomz-7b1 .273 ± .016 .483 ± .014 .001 ± .000
EuroLLM-1.7B .255 ± .005 .445 ± .026 .033 ± .004
EuroLLM-1.7B-Instruct .264 ± .009 .469 ± .037 .083 ± .006
Gemma-1.1-7b-Instruct .407 ± .010 .458 ± .029 .117 ± .020
Gemma-2-27b-Instruct .672 ± .012 .611 ± .016 .749 ± .019
Gemma-2-9b-Instruct .577 ± .009 .587 ± .004 .428 ± .021
Gemma-7b .558 ± .015 .495 ± .030 .391 ± .025
Meta-Llama-2-13B-Chat .416 ± .014 .474 ± .046 .147 ± .030
Meta-Llama-2-7B .346 ± .040 .450 ± .033 .074 ± .010
Meta-Llama-2-7B-Chat .355 ± .022 .485 ± .038 .093 ± .028
Meta-Llama-3-8B .531 ± .015 .485 ± .015 .323 ± .027
Meta-Llama-3-8B-Instruct .524 ± .019 .525 ± .024 .536 ± .040
Meta-Llama-3.1-70B-Instruct .764 ± .020 .574 ± .017 .666 ± .088
Meta-Llama-3.1-8B .534 ± .017 .502 ± .021 .346 ± .022
Meta-Llama-3.1-8B-Instruct .554 ± .015 .532 ± .016 .543 ± .039
Mistral-7B-Instruct-v0.1 .418 ± .017 .516 ± .045 .171 ± .029
Mistral-7B-Instruct-v0.2 .474 ± .018 .589 ± .051 .258 ± .023
Mistral-7B-Instruct-v0.3 .493 ± .013 .548 ± .046 .320 ± .019
Mistral-7B-v0.1 .497 ± .021 .495 ± .046 .250 ± .026
Mistral-7B-v0.3 .498 ± .016 .486 ± .042 .238 ± .034
Mistral-NeMo-Minitron-8B-Base .526 ± .032 .481 ± .034 .384 ± .028
Mistral-Nemo-Base-12.2B_2407 .585 ± .011 .524 ± .026 .422 ± .034
Mistral-Nemo-Instruct-12.2B_2407 .579 ± .009 .588 ± .023 .549 ± .025
Mixtral-8x7B-Instruct-v0.1 .609 ± .013 .620 ± .045 .488 ± .027
Mixtral-8x7B-v0.1 .615 ± .017 .505 ± .033 .429 ± .029
Occiglot-7b-eu5 .394 ± .036 .451 ± .040 .154 ± .018
Occiglot-7b-eu5-Instruct .402 ± .027 .465 ± .051 .194 ± .030
Pharia-1-LLM-7B-control .291 ± .014 .448 ± .037 .010 ± .005
Pharia-1-LLM-7B-control-aligned .297 ± .024 .447 ± .040 .013 ± .009
Phi-3-medium-14B-128k-Instruct .516 ± .016 .472 ± .032 .432 ± .042
Phi-3-medium-14B-4k-Instruct .526 ± .022 .470 ± .036 .470 ± .048
Phi-3-mini-3.8B-128k-Instruct .387 ± .033 .485 ± .036 .193 ± .045
Phi-3-mini-4k-Instruct .376 ± .033 .476 ± .038 .165 ± .048
Qwen2-7B .578 ± .028 .534 ± .021 .584 ± .058
Qwen2-7B-Instruct .572 ± .022 .547 ± .024 .512 ± .064
Vicuna-13b-v1.5 .437 ± .014 .509 ± .039 .182 ± .025
Vicuna-33b-v1.3 .465 ± .019 .529 ± .047 .149 ± .033
c4ai-command-r-35B-v01 .540 ± .045 .542 ± .021 .424 ± .053
Model ARC GSM8K HellaSwag MMLU TruthfulQA
Aya-23-8B .216 .155 .230 .109 .006
Bloom-7b1 .126 .006 .139 .006 -0.047
Bloomz-7b1 .128 -$0$141 .056 -0.025
EuroLLM-1.7B .052 .004 .050 -0.001 -0.031
EuroLLM-1.7B-Instruct .056 .023 .052 .003 -0.030
Gemma-1.1-7b-Instruct .094 .047 .107 .061 .003
Gemma-2-27b-Instruct .042 .032 .058 .035 .015
Gemma-2-9b-Instruct .084 .061 .094 .055 .023
Gemma-7b .076 .047 .084 .043 -0.008
Meta-Llama-2-13B-Chat .152 .120 .160 .082 .002
Meta-Llama-2-7B .146 .042 .144 .071 -0.021
Meta-Llama-2-7B-Chat .149 .073 .146 .073 .005
Meta-Llama-3-8B .107 .076 .116 .069 -0.023
Meta-Llama-3-8B-Instruct .116 .116 .123 .079 .012
Meta-Llama-3.1-70B-Instruct .055 .029 .076 .030 .009
Meta-Llama-3.1-8B .104 .073 .122 .065 -0.012
Meta-Llama-3.1-8B-Instruct .108 .096 .122 .071 .013
Mistral-7B-Instruct-v0.1 .172 .133 .156 .097 .057
Mistral-7B-Instruct-v0.2 .177 .119 .179 .096 .073
Mistral-7B-Instruct-v0.3 .177 .126 .171 .096 .044
Mistral-7B-v0.1 .169 .106 .165 .101 -0.008
Mistral-7B-v0.3 .170 .095 .164 .097 -0.004
Mistral-NeMo-Minitron-8B-Base .175 .111 .176 .103 .012
Mistral-Nemo-Base-12.2B_2407 .108 .087 .138 .070 .001
Mistral-Nemo-Instruct-12.2B_2407 .108 .107 .138 .069 -0.005
Mixtral-8x7B-Instruct-v0.1 .171 .140 .183 .103 .053
Mixtral-8x7B-v0.1 .165 .128 .172 .103 .007
Occiglot-7b-eu5 .197 .100 .219 .112 -0.028
Occiglot-7b-eu5-Instruct .202 -$0$224 .114 -0.016
Pharia-1-LLM-7B-control .276 .045 .288 .124 .000
Pharia-1-LLM-7B-control-aligned .281 .104 .294 .136 .027
Phi-3-medium-14B-128k-Instruct .259 .320 .256 .188 .064
Phi-3-medium-14B-4k-Instruct .262 .317 .260 .193 .072
Phi-3-mini-3.8B-128k-Instruct .248 .390 .228 .195 .059
Phi-3-mini-4k-Instruct .245 .363 .233 .183 .070
Qwen2-7B .182 .156 .181 .099 .016
Qwen2-7B-Instruct .184 .082 .188 .098 .038
Vicuna-13b-v1.5 .148 .092 .161 .088 .015
Vicuna-33b-v1.3 .182 .117 .174 .101 .024
c4ai-command-r-35B-v01 .201 .124 .205 .112 .030
Table 16: Differences between high and medium resource languages
Model Name PC
Aya-23-8B 0.571
Meta-Llama-3-8B-Instruct 0.562
Mistral-7B-Instruct-v0.3 0.562
Meta-Llama-3-8B 0.556
gemma-7b 0.555
Llama-2-7b-chat-hf 0.553
Meta-Llama-3.1-8B 0.552
gemma-1.1-7b-it 0.552
Meta-Llama-3.1-8B-Instruct 0.549
occiglot-7b-eu5 0.548
occiglot-7b-eu5-instruct 0.546
Mistral-7B-v0.1 0.542
Mistral-7B-Instruct-v0.2 0.539
Mistral-7B-v0.3 0.536
Phi-3-mini-4k-instruct 0.534
Mistral-7B-Instruct-v0.1 0.528
Llama-2-7b-hf 0.520
bloomz-7b1 0.439
bloom-7b1 0.401
Table 17: Model Performance Correlation Matrix (Sorted by Correlation Coefficient in Descending Order)
(PC: Pearson Correleation Coefficient)
Language PC
es 0.589991
fr 0.586389
it 0.567793
sv 0.564892
da 0.557272
nl 0.555682
ro 0.546222
de 0.545930
sk 0.473632
hu 0.444939
Table 18: Language Correlation Coefficients (Sorted by Coefficient in Descending Order)
(PC: Pearson Correleation Coefficient)
Task PC
EU21-ARC 0.610
EU21-HeSw 0.543
EU21-MMLU 0.538
Table 19: Task Correlation Coefficients (Sorted by Coefficient in Descending Order)
(PC: Pearson Correleation Coefficient)
Figure 5: Task Accuracies per Language for EU21-ARC.
Figure 6: Task Accuracies per Language for EU21-HeSw.
Figure 7: Task Accuracies per Language for EU21-MMLU.
Figure 8: Task Accuracies per Language for EU21-TQA.
Figure 9: Task Accuracies per Language for EU21-GSM8k.
Appendix C Correlation with Human Preference Scores
Language Number of Battles
English 1,035,743
German 50,419
Spanish 31,278
French 28,819
Portuguese 16,959
Italian 14,412
Czech 7,517
Polish 6,868
Dutch 2,748
Danish 2,392
Swedish 2,325
Finnish 2,211
Hungarian 2,066
Bulgarian 1,737
Romanian 1,273
Greek 968
Slovak 831
Croatian 411
Lithuanian 395
Estonian 256
Slovenian 232
Latvian 199
Irish 134
Maltese 91
Table 20: Battle Overview
Model and Hugging Face Link
c4ai-command-r-35b-v01
gemma-1.1-7b-it
gemma-2-27b-it
gemma-2-9b-Instruct
Llama-2-7b-chat-hf
Llama-2-13b-chat-hf
Meta-Llama-3-8B-Instruct
Meta-Llama-3.1-8B-Instruct
Meta-Llama-3.1-70B-Instruct
Mistral-7B-Instruct-v0.1
Mistral-7B-Instruct-v0.2
Mixtral-8x7B-Instruct-v0.1
Phi-3-medium-14b-4k-instruct
Phi-3-mini-128k-Instruct
Phi-3-mini-4k-Instruct
Vicuna-13b-v1.5
Vicuna-33b-v1.3
Table 21: Model Overview
Table 22: Correlation Coefficients for Task EU20-ARC between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Task Lang. Pearson Pearson CI Spearman N NUR
EU20-ARC CS .600@.00541(+).600@.00541.600@.$0$00541 ( + ) (.$(.249,1)$$499@.02070\,(+).$ 17 $9\,(+)$
EU20-ARC EN .$722@.00054\,(+).$ (.$(.439,1)$$691@.00106\,(+).$ 17 $16\,(+)$
EU20-ARC FR .$772@.00014\,(+).$ (.$(.526,1)$$770@.00015\,(+).$ 17 $14\,(+)$
EU20-ARC DE .$815@.00003\,(+).$ (.$(.606,1)$$779@.00011\,(+).$ 17 $14\,(+)$
EU20-ARC IT .700@.00088(+).700@.00088.700@.$0$00088 ( + ) (.$(.403,1)$$696@.00096\,(+).$ 17 $15\,(+)$
EU20-ARC PL .$868@.00000\,(+).$ (.$(.709,1)$$836@.00001\,(+).$ 17 $13\,(+)$
EU20-ARC PT .$779@.00012\,(+).$ (.$(.539,1)$$784@.00010\,(+).$ 17 $14\,(+)$
EU20-ARC ES .$743@.00032\,(+).$ (.$(.475,1)$$760@.00020\,(+).$ 17 $14\,(+)$
Table 22: Correlation Coefficients for Task EU20-ARC between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Table 23: Correlation Coefficients for Task EU20-GSM8K between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Task Lang. Pearson Pearson CI Spearman N NUR
EU20-GSM8K CS .$465@.03006\,(+).$ (.$(.064,1)$$302@.11966\,(-).$ 17 $9\,(+)$
EU20-GSM8K EN .$671@.00159\,(+).$ (.$(.357,1)$$603@.00520\,(+).$ 17 $16\,(+)$
EU20-GSM8K FR .$723@.00052\,(+).$ (.$(.441,1)$$694@.00101\,(+).$ 17 $14\,(+)$
EU20-GSM8K DE .$717@.00060\,(+).$ (.$(.431,1)$$613@.00441\,(+).$ 17 $14\,(+)$
EU20-GSM8K IT .$616@.00422\,(+).$ (.$(.272,1)$$586@.00674\,(+).$ 17 $15\,(+)$
EU20-GSM8K PL .$785@.00010\,(+).$ (.$(.549,1)$$713@.00065\,(+).$ 17 $13\,(+)$
EU20-GSM8K PT .$786@.00009\,(+).$ (.$(.553,1)$$743@.00032\,(+).$ 17 $14\,(+)$
EU20-GSM8K ES .$685@.00119\,(+).$ (.$(.380,1)$$662@.00190\,(+).$ 17 $14\,(+)$
Table 23: Correlation Coefficients for Task EU20-GSM8K between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Table 24: Correlation Coefficients for Task EU20-HellaSwag between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Task Lang. Pearson Pearson CI Spearman N NUR
EU20-HellaSwag CS .$581@.00720\,(+).$ (.$(.221,1)$$479@.02574\,(+).$ 17 $9\,(+)$
EU20-HellaSwag EN .$430@.04248\,(+).$ (.$(.020,1)$$436@.03999\,(+).$ 17 $16\,(+)$
EU20-HellaSwag FR .$722@.00053\,(+).$ (.$(.440,1)$$679@.00136\,(+).$ 17 $14\,(+)$
EU20-HellaSwag DE .$719@.00057\,(+).$ (.$(.435,1)$$762@.00019\,(+).$ 17 $14\,(+)$
EU20-HellaSwag IT .$714@.00064\,(+).$ (.$(.426,1)$$669@.00165\,(+).$ 17 $15\,(+)$
EU20-HellaSwag PL .$851@.00001\,(+).$ (.$(.676,1)$$855@.00001\,(+).$ 17 $13\,(+)$
EU20-HellaSwag PT .$729@.00045\,(+).$ (.$(.452,1)$$784@.00010\,(+).$ 17 $14\,(+)$
EU20-HellaSwag ES .$687@.00117\,(+).$ (.$(.381,1)$$752@.00025\,(+).$ 17 $14\,(+)$
Table 24: Correlation Coefficients for Task EU20-HellaSwag between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Table 25: Correlation Coefficients for Task EU20-MMLU between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Task Lang. Pearson Pearson CI Spearman N NUR
EU20-MMLU CS .$564@.00917\,(+).$ (.$(.197,1)$$416@.04851\,(-).$ 17 $9\,(+)$
EU20-MMLU EN .$737@.00037\,(+).$ (.$(.465,1)$$610@.00464\,(+).$ 17 $16\,(+)$
EU20-MMLU FR .$814@.00004\,(+).$ (.$(.603,1)$$804@.00005\,(+).$ 17 $14\,(+)$
EU20-MMLU DE .$792@.00008\,(+).$ (.$(.563,1)$$782@.00010\,(+).$ 17 $14\,(+)$
EU20-MMLU IT .$734@.00040\,(+).$ (.$(.460,1)$$610@.00464\,(+).$ 17 $15\,(+)$
EU20-MMLU PL .$885@.00000\,(+).$ (.$(.744,1)$$863@.00000\,(+).$ 17 $13\,(+)$
EU20-MMLU PT .$804@.00005\,(+).$ (.$(.585,1)$$760@.00020\,(+).$ 17 $14\,(+)$
EU20-MMLU ES .$768@.00016\,(+).$ (.$(.520,1)$$770@.00015\,(+).$ 17 $14\,(+)$
Table 25: Correlation Coefficients for Task EU20-MMLU between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Table 26: Correlation Coefficients for Task EU20-TruthfulQA between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Task Lang. Pearson Pearson CI Spearman N NUR
EU20-TruthfulQA CS .$433@.04129\,(+).$ (.$(.024,1)$$401@.05534\,(-).$ 17 $9\,(+)$
EU20-TruthfulQA EN .$426@.04397\,(+).$ (.$(.016,1)$$510@.01828\,(+).$ 17 $16\,(+)$
EU20-TruthfulQA FR .$360@.07793\,(-).$ $(-0.063,1)$$309@.11389\,(-).$ 17 $14\,(+)$
EU20-TruthfulQA DE .$343@.08872\,(-).$ $(-0.082,1)$$417@.04808\,(-).$ 17 $14\,(+)$
EU20-TruthfulQA IT .$380@.06619\,(-).$ $(-0.039,1)$$277@.14092\,(-).$ 17 $15\,(+)$
EU20-TruthfulQA PL .$603@.00524\,(+).$ (.$(.252,1)$$706@.00077\,(+).$ 17 $13\,(+)$
EU20-TruthfulQA PT .$548@.01138\,(+).$ (.$(.174,1)$$569@.00861\,(+).$ 17 $14\,(+)$
EU20-TruthfulQA ES .$447@.03615\,(+).$ (.$(.041,1)$$473@.02757\,(+).$ 17 $14\,(+)$
Table 26: Correlation Coefficients for Task EU20-TruthfulQA between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Table 27: Correlation Coefficients for Task Avg. between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Task Lang. Pearson Pearson CI Spearman N NUR
Avg. CS .$570@.00847\,(+).$ (.$(.205,1)$$430@.04230\,(+).$ 17 $9\,(+)$
Avg. EN .$716@.00062\,(+).$ (.$(.429,1)$$618@.00412\,(+).$ 17 $16\,(+)$
Avg. FR .$770@.00015\,(+).$ (.$(.523,1)$$745@.00030\,(+).$ 17 $14\,(+)$
Avg. DE .$775@.00013\,(+).$ (.$(.531,1)$$740@.00034\,(+).$ 17 $14\,(+)$
Avg. IT .$694@.00101\,(+).$ (.$(.393,1)$$674@.00150\,(+).$ 17 $15\,(+)$
Avg. PL .$895@.00000\,(+).$ (.$(.765,1)$$858@.00001\,(+).$ 17 $13\,(+)$
Avg. PT .$811@.00004\,(+).$ (.$(.598,1)$$806@.00005\,(+).$ 17 $14\,(+)$
Avg. ES .$751@.00025\,(+).$ (.$(.490,1)$$784@.00010\,(+).$ 17 $14\,(+)$
Table 27: Correlation Coefficients for Task Avg. between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Table 28: Correlation Coefficients for Task Reasoning Avg. between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Task Lang. Pearson Pearson CI Spearman N NUR
Avg - TQA. CS .$572@.00818\,(+).$ (.$(.208,1)$$397@.05715\,(-).$ 17 $9\,(+)$
Avg - TQA. EN .$722@.00053\,(+).$ (.$(.440,1)$$623@.00380\,(+).$ 17 $16\,(+)$
Avg - TQA. FR .$793@.00007\,(+).$ (.$(.564,1)$$789@.00008\,(+).$ 17 $14\,(+)$
Avg - TQA. DE .$792@.00008\,(+).$ (.$(.562,1)$$775@.00013\,(+).$ 17 $14\,(+)$
Avg - TQA. IT .$707@.00076\,(+).$ (.$(.415,1)$$664@.00182\,(+).$ 17 $15\,(+)$
Avg - TQA. PL .$901@.00000\,(+).$ (.$(.777,1)$$855@.00001\,(+).$ 17 $13\,(+)$
Avg - TQA. PT .$819@.00003\,(+).$ (.$(.613,1)$$7$00012 ( + ) 17 $14\,(+)$
Avg - TQA. ES .$756@.00022\,(+).$ (.$(.498,1)$$767@.00016\,(+).$ 17 $14\,(+)$
Table 28: Correlation Coefficients for Task Reasoning Avg. between Few-Shot Accuracy and LMSYS ELO Score, incl. 95% confidence intervals for Pearson. Numbers behind @ are p values, $+$ indicates value below 0.05. NUR: num. unique ranks with  95% confidence, $+$ indicates value below 50%.
Appendix D Influence of translation service
Figure 10: Accuracy comparison of different models on our and OKAPIs translated ARC dataset.
Figure 11: Accuracy comparison of different models on our and OKAPIs translated HELLASWAG dataset.
Figure 12: Accuracy comparison of different models on our and OKAPIs translated MMLU dataset.
Table 29: Comparison of ogx and okapi scores for ARC and Hellaswag benchmarks (Higher values in bold)
Language ARC Hellaswag MMLU
OGX OKAPI OGX OKAPI OGX OKAPI
German 0.8289 0.8289 0.7857 0.7971 0.8110 0.8098
Spanish 0.8509 0.8503 0.7919 0.8050 0.8295 0.8287
French 0.8479 0.8471 0.8032 0.8122 0.8269 0.8236
Italian 0.8532 0.8516 0.8014 0.8133 0.8297 0.8261
Hungarian 0.8489 0.8406 0.8041 0.7896 0.8249 0.8135
Danish 0.8369 0.8321 0.8024 0.8058 0.8150 0.8106
Romanian 0.8515 0.8487 0.8068 0.7440 0.8336 0.8273
Swedish 0.8465 0.8426 0.8157 0.8155 0.8249 0.8212
Slovak 0.8493 0.8384 0.7982 0.7788 0.8238 0.8102
Dutch 0.8364 - 0.8002 - 0.8177 -
Portuguese 0.8386 - 0.7755 - 0.8233 -
Slovenian 0.8424 - 0.7969 - 0.8258 -
Polish 0.8367 - 0.7807 - 0.8090 -
Latvian 0.8268 - 0.7846 - 0.8083 -
Lithuanian 0.8324 - 0.7896 - 0.8123 -
Finnish 0.8510 - 0.8197 - 0.8333 -
Estonian 0.8275 - 0.8044 - 0.8062 -
Bulgarian 0.8602 - 0.8114 - 0.8350 -
Czech 0.8456 - 0.7920 - 0.8178 -
Table 30: Average sample number for each task of eleven translations.
Split Original OGX OKAPI
MMLU Val. 1531 276.18 ±$\pm 8$
Test 14042 13198.09 ±$\pm 101$
Hellaswag Val 10042 10035.18 ±$\pm 19$ 9276.63 ±$\pm 112$
ARC Val. 299 299 272.66 ±$\pm 0.5$
Test 1172 1172 1071 ±$\pm 2$
TruthfulQA Val. 817 817 782.09 ±$\pm 6$
Table 31: Hellaswag translation examples compared between OGX and OKAPI
Lang./Dataset Prompt
Score
English/Original [header] How to choose a good budgie [title] Locate budgies for sale from a reputable seller. [step] You can find budgies for sale online, in want-ads, or in pet stores. You can also get one from your local animal shelter. - Most animal shelters sell budgies for a flat fee, so this will limit your bank expenses. [substeps] You might also be able to compare prices to get a good deal. - You’ll need to get a price point that is reasonable. Here’s a guide to locate an experienced seller : [substeps] Generally , 20-50% off the appraised price is reasonable. - Wherever you get your budgie from, make sure the seller has healthy and humanely treated birds. [substeps] Look at online reviews for the sellers you find. - Basically, you need to find a good seller who has had previous owners. [substeps] Check to see if the parent owned the budgie.
German/OGX COMET: 0.6329 Wie man einen guten Wellensittich auswählt. Suchen Sie nach Wellensittichen, die von einem seriösen Verkäufer verkauft werden. Sie können Wellensittiche im Internet, in Kleinanzeigen oder in Zoohandlungen finden, die zum Verkauf stehen. Sie können auch einen Wellensittich aus Ihrem örtlichen Tierheim bekommen. - Die meisten Tierheime verkaufen Wellensittiche gegen eine Pauschalgebühr, so dass sich Ihre Bankkosten in Grenzen halten. Vielleicht können Sie auch die Preise vergleichen, um ein gutes Angebot zu bekommen. - Sie müssen einen angemessenen Preis finden. Im Allgemeinen sind 20-50 % des Schätzpreises angemessen. - Wo auch immer Sie Ihren Wellensittich kaufen, stellen Sie sicher, dass der Verkäufer gesunde und artgerecht gehaltene Vögel hat. Schauen Sie sich die Online-Bewertungen der Verkäufer an, die Sie finden. - Grundsätzlich müssen Sie einen guten Verkäufer finden, der bereits Vorbesitzer hatte. Erkundigen Sie sich, ob der Wellensittich den Eltern gehört hat.
German/OKAPI COMET: 0.6983 [header] Wie man einen guten Wellensittich auswählt [title] Lokalisieren Sie Wellensittiche zum Verkauf von einem seriösen Verkäufer. [step] Sie können Wellensittiche zum Verkauf online, in Kleinanzeigen oder in Zoohandlungen finden. Sie können auch einen aus Ihrem örtlichen Tierheim bekommen. - Die meisten Tierheime verkaufen Wellensittiche zu einem Festpreis, was Ihre Bankkosten begrenzen wird. [substeps] Sie können auch Preise vergleichen, um ein gutes Angebot zu erhalten. - Sie müssen einen vernünftigen Preis finden. Hier ist ein Leitfaden, um einen erfahrenen Verkäufer zu finden: [substeps] In der Regel sind 20-50% Rabatt auf den geschätzten Preis angemessen. - Wo immer Sie Ihren Wellensittich her bekommen, stellen Sie sicher, dass der Verkäufer gesunde und menschlich behandelte Vögel hat. [substeps] Schauen Sie sich Online-Bewertungen für die Verkäufer an, die Sie finden. - Im Grunde genommen müssen Sie einen guten Verkäufer finden, der bereits Vorbesitzer hatte. [substeps] Überprüfen Sie, ob der Vorbesitzer den Wellensittich besaß.
English/Original A green title page appears, with words showing that the video will be of a tenor saxophone player as well as the website to receive to lessons. A black male then appears with a green sweater and black hat playing the sax. the man - bluffs on the drums and the video continues, at various points, to show individuals playing the instrument. - continues to play and cheers, shakes hands with the man, throws a napkin, writes on the napkin and gives him a card. - then lays the instrument across his lap and begins talking about what he is doing. - jumps into a black and white video and temporarily stops playing, saying, " big calibers ".
French/OGX COMET: 0.5770 Une page de titre verte apparaît, avec des mots indiquant que la vidéo sera celle d’un joueur de saxophone ténor, ainsi que le site web permettant de recevoir des leçons. Un homme noir apparaît ensuite, avec un pull vert et un chapeau noir, jouant du saxophone. l’homme - bluffe à la batterie et la vidéo continue, à différents moments, à montrer des personnes jouant de l’instrument. - - continue à jouer et applaudit, serre la main de l’homme, lui lance une serviette, écrit sur la serviette et lui donne une carte. pose ensuite l’instrument sur ses genoux et commence à parler de ce qu’il fait. - saute dans une vidéo en noir et blanc et s’arrête temporairement de jouer en disant " big calibers " (gros calibres).
French/OKAPI COMET: 0.7639 Une page de titre verte apparaît, avec des mots indiquant que la vidéo sera d’un joueur de saxophone ténor ainsi que le site Web pour recevoir des leçons. Un homme noir apparaît alors avec un pull vert et un chapeau noir en jouant du saxophone. l’homme - bluffe à la batterie et la vidéo continue, à différents moments, de montrer des individus jouant de l’instrument. - continue à jouer et crier, serre la main de l’homme, jette une serviette, écrit sur la serviette et lui donne une carte. - puis pose l’instrument sur ses genoux et commence à parler de ce qu’il fait. - saut dans une vidéo en noir et blanc et arrête temporairement de jouer, en disant, "gros calibres".
English/Original A man drinks from a can next to a hole in the snow. he - swims slowly down a shallow body of water. - speaks and gestures to the camera. - gets his teeth bitten by a young man. - is drawn in with pick and shovel.
Spanish/OGX COMET: 0.7803 Un hombre bebe de una lata junto a un agujero en la nieve. él - nada lentamente por una masa de agua poco profunda. - habla y gesticula a la cámara. - un joven le muerde los dientes. - se acerca con pico y pala.
Spanish/OKAPI COMET: 0.7082 Un hombre bebe de una lata junto a un agujero en la nieve. él - nada lentamente por un cuerpo de agua poco profundo. - habla y gestiona hacia la cámara. - recibe una mordida de un joven en los dientes. - es sacado con pico y pala.
English/Original [header] How to ease a stomach virus [title] Avoid solid foods for a few hours to let your stomach settle. [step] When you first notice symptoms of the stomach virus, you should allow your stomach to settle. Eating food early on could irritate the stomach lining more, worsening your symptoms. - For a few hours, avoid solid foods. [title] Take small sips of water or chew ice chips. - [title] Avoid eating something for at least 10 minutes. [step] Studies have not found that eating for a few hours can relieve stomach symptoms. - Trying to eat a solid meal a few hours before your nausea attack will only make symptoms worse. If you feel ill after the onset of the symptoms, stop eating. - [substeps] Try to eat a few hours before work or school, particularly if you’re feeling sick. Moving around a bit can also minimize your nausea and help your stomach to settle.
Italian/OGX COMET: 0.8592 Come alleviare un virus intestinale. Evitate i cibi solidi per qualche ora per permettere allo stomaco di calmarsi. Quando si avvertono i primi sintomi di un virus intestinale, è necessario lasciare che lo stomaco si stabilizzi. Mangiare cibo subito potrebbe irritare maggiormente la mucosa gastrica, peggiorando i sintomi. - Per qualche ora, evitate i cibi solidi. Bevete piccoli sorsi d’acqua o masticate del ghiaccio. - Evitare di mangiare qualcosa per almeno 10 minuti. Gli studi non hanno riscontrato che mangiare per qualche ora possa alleviare i sintomi dello stomaco. - Cercare di mangiare un pasto solido qualche ora prima dell’attacco di nausea non farà altro che peggiorare i sintomi. Se vi sentite male dopo l’inizio dei sintomi, smettete di mangiare. - Cercate di mangiare qualche ora prima del lavoro o della scuola, soprattutto se vi sentite male. Anche muoversi un po’ può ridurre la nausea e aiutare lo stomaco a stabilizzarsi.
Italian/OKAPI COMET: 0.6923 [header] Come alleviare un virus dello stomaco [title] Evita i cibi solidi per alcune ore per far riposare lo stomaco. [step] Quando noti i primi sintomi del virus dello stomaco, dovresti far riposare lo stomaco. Mangiare subito potrebbe irritare ancora di più la mucosa gastrica, peggiorando i sintomi. - Per alcune ore, evita i cibi solidi. [title] Fai piccoli sorsi d’acqua o mastica cubetti di ghiaccio. - [title] Evita di mangiare qualcosa per almeno 10 minuti. [step] Gli studi non hanno trovato alcuna relazione tra il mangiare e il sollievo dei sintomi gastrici. - Prova a mangiare qualche ora prima del lavoro o della scuola, in particolare se ti senti male. Muoverti un po’ può anche ridurre la nausea e aiutare lo stomaco a riposare. - Cercare di mangiare un pasto solido qualche ora prima dell’attacco di nausea renderà solo i sintomi peggiori. Se ti senti male dopo l’inizio dei sintomi, smetti di mangiare.
English/Original A man wearing headphones is in a studio, speaking into a microphone. Other men are shown, one of which begins playing a harmonica. he - stops playing and says something that makes the dj laugh. - is seen playing intensely with the camera. - plays with a jack o lantern and leaves the room. - seems very pleased with the information he has gathered as he allows the host to speak.
Swedish/OGX COMET: 0.8174 En man med hörlurar befinner sig i en studio och talar i en mikrofon. Andra män visas, varav en börjar spela munspel. han - slutar spela och säger något som får dj:n att skratta. - ses leka intensivt med kameran. - leker med en jack o lantern och lämnar rummet. - verkar mycket nöjd med den information han har samlat in när han låter programledaren tala.
Swedish/OKAPI COMET: 0.7227 En man som bär hörlurar är i en studio och talar i en mikrofon. Andra män visas, varav en börjar spela munspel. han - slutar spela och säger något som får DJ:n att skratta. - ses spela intensivt med kameran. - spelar med en Halloween-lampa och lämnar rummet. - verkar mycket nöjd med informationen han har samlat in och låter värdet tala.
English/Original [header] How to officiate at a nondenominational funeral service [title] Talk to the family. [step] Call the family and arrange to meet them, preferably in their home. Meeting them in public is ok in a bind, but if that’s the arrangement, try not to meet more than two to three people. - [substeps] If the ceremony is inside a church, don’t bother sitting outside on a park bench. If you’re a minister or midwife, it’s best to arrive by the usual time of worship. - [substeps] Of course, only with a family member and one of their kids is okay. If no one at your funeral is willing to attend, that’s fine. - People aren’t always happy with the small acts of an officiant. In some cases, the family can become drunk, agitated, or violent during or around a service. - If a funeral home is taking care of other arrangements, the funeral directors may have space at the funeral home where you could meet. Invite the family to have other family members and close friends present if they wish, but be sensitive to their needs and to the space considerations.
Hungarian/OGX COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni. Beszélj a családdal. Hívja fel a családot, és egyeztessen velük egy találkozót, lehetőleg az otthonukban. A nyilvános találkozás kötöttségek esetén rendben van, de ha ez a megállapodás, próbáljon meg nem két-három embernél többel találkozni. - Ha a szertartás templomban lesz, ne fáradjon azzal, hogy kint üljön egy parkban lévő padon. Ha lelkész vagy bába, a legjobb, ha a szokásos istentiszteleti időre érkezik. - Természetesen csak egy családtaggal és az egyik gyermekükkel együtt nem baj. Ha a temetéseden senki sem hajlandó részt venni, az sem baj. - Az emberek nem mindig örülnek a szertartásvezető apró cselekedeteinek. Egyes esetekben a család részeggé, izgatottá vagy erőszakossá válhat a szertartás alatt vagy körül. - Ha egy temetkezési vállalat gondoskodik egyéb előkészületekről, a temetkezési vállalatnak lehet, hogy van olyan hely a ravatalozóban, ahol találkozhatnátok. Hívja meg a családot, hogy más családtagok és közeli barátok is jelen legyenek, ha kívánják, de legyen tekintettel az ő igényeikre és a térbeli megfontolásokra.
Hungarian/OKAPI COMET: 0.7061 [header] Hogyan végezzen szertartást nem vallásos temetésen [title] Beszéljen a családdal. [step] Hívja fel a családot és beszéljen meg egy találkozót, lehetőleg az otthonukban. Nyilvános helyen találkozni is lehet, ha nincs más lehetőség, de ha így van, akkor ne találkozzon több mint két-három emberrel. - [substeps] Ha a szertartást templomban tartják, ne üljön le kinn egy parki padra. Ha lelkész vagy bábák, a szokásos istentiszteleti időben érkezzen. - [substeps] Természetesen csak egy családtaggal és az egyik gyerekkel találkozni is lehet. Ha senki sem akar az Ön temetésén részt venni, az sem baj. - Az emberek nem mindig elégedettek a szertartás megtartásával, még akkor sem, ha az pontosan el lett végezve. Néhány esetben a család részére az esemény körüli korlátozások miatt ajtóstul jön a stressz, ezt jogosak. - Ha a temetkezési vállalkozás többi rendezvényét szervezi, a temetkezés szervezőinek lehetősége nyílhat arra, hogy találkozzanak a temetőnél. Azonban ha valaki akarja, hogy ott legyenek családtagok vagy közel álló barátok is, tartsuk észben a helyzetet és az igényeket.
English/Original A man is running while holding a pole on a track. he - uses the pole to vault through the air. - runs down the track, wrapping his legs around the pole. - runs into a sand pit and starts to jump. - lands on a pile of sand.
Slovak/OGX COMET: 0.8336 Muž beží a pritom drží tyč na dráhe. on - používa tyč na skok vo vzduchu. - beží po dráhe a obopína nohy okolo tyče. - vbehne do pieskoviska a začne skákať. - dopadne na hromadu piesku.
Slovak/OKAPI COMET: 0.7494 Muž beží po dráhe a drží tyč. on - používa tyč na skákanie cez vzduch. - beží po dráhe a obopína nohy okolo tyče. - beží do pieskového výskoku. - pristane na kupke piesku.
English/Original A small girl runs away from the pinata while holding the stick. The child returns and hits the pinata one time. the people - continue to remove key ingredients from the bags. - cheer for the child getting hit. - watch while the girl throws the stick twice. - all clap and the girl smiles and turns toward the camera.
Danish/OGX COMET: 0.8450 En lille pige løber væk fra pinataen, mens hun holder pinden. Barnet vender tilbage og rammer pinataen én gang. personerne - fortsætter med at fjerne nøgleingredienser fra poserne. - jubler over barnet, der bliver ramt. - ser til, mens pigen kaster pinden to gange. - alle klapper, og pigen smiler og vender sig mod kameraet.
Danish/OKAPI COMET: 0.7696 En lille pige løber væk fra piñataen, mens hun holder stokken. Barnet vender tilbage og rammer piñataen en gang. folkene - fortsætter med at fjerne vigtige ingredienser fra poserne. - hepper på, at barnet rammer piñataen. - ser mens pigen kaster stokken to gange. - klapper alle og pigen smiler og vender sig mod kameraet.
English/Original A man is sitting on a roof. he - is using wrap to wrap a pair of skis. - is ripping level tiles off. - is holding a rubik’s cube. - starts pulling up roofing on a roof.
[…]
English/Original [header] How to officiate at a nondenominational funeral service [title] Talk to the family. [step] Call the family and arrange to meet them, preferably in their home. Meeting them in public is ok in a bind, but if that’s the arrangement, try not to meet more than two to three people. - [substeps] If the ceremony is inside a church, don’t bother sitting outside on a park bench. If you’re a minister or midwife, it’s best to arrive by the usual time of worship. - [substeps] Of course, only with a family member and one of their kids is okay. If no one at your funeral is willing to attend, that’s fine. - People aren’t always happy with the small acts of an officiant. In some cases, the family can become drunk, agitated, or violent during or around a service. - If a funeral home is taking care of other arrangements, the funeral directors may have space at the funeral home where you could meet. Invite the family to have other family members and close friends present if they wish, but be sensitive to their needs and to the space considerations.
Hungarian/OGX COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni. Beszélj a családdal. Hívja fel a családot, és egyeztessen velük egy találkozót, lehetőleg az otthonukban. A nyilvános találkozás kötöttségek esetén rendben van, de ha ez a megállapodás, próbáljon meg nem két-három embernél többel találkozni. - Ha a szertartás templomban lesz, ne fáradjon azzal, hogy kint üljön egy parkban lévő padon. Ha lelkész vagy bába, a legjobb, ha a szokásos istentiszteleti időre érkezik. - Természetesen csak egy családtaggal és az egyik gyermekükkel együtt nem baj. Ha a temetéseden senki sem hajlandó részt venni, az sem baj. - Az emberek nem mindig örülnek a szertartásvezető apró cselekedeteinek. Egyes esetekben a család részeggé, izgatottá vagy erőszakossá válhat a szertartás alatt vagy körül. - Ha egy temetkezési vállalat gondoskodik egyéb előkészületekről, a temetkezési vállalatnak lehet, hogy van olyan hely a ravatalozóban, ahol találkozhatnátok. Hívja meg a családot, hogy más családtagok és közeli barátok is jelen legyenek, ha kívánják, de legyen tekintettel az ő igényeikre és a térbeli megfontolásokra.
Hungarian/OKAPI COMET: 0.7060 [header] Hogyan végezzen szertartást nem vallásos temetésen [title] Beszéljen a családdal. [step] Hívja fel a családot és beszéljen meg egy találkozót, lehetőleg az otthonukban. Nyilvános helyen találkozni is lehet, ha nincs más lehetőség, de ha így van, akkor ne találkozzon több mint két-három emberrel. - [substeps] Ha a szertartást templomban tartják, ne üljön le kinn egy parki padra. Ha lelkész vagy bábák, a szokásos istentiszteleti időben érkezzen. - [substeps] Természetesen csak egy családtaggal és az egyik gyerekkel találkozni is lehet. Ha senki sem akar az Ön temetésén részt venni, az sem baj. - Az emberek nem mindig elégedettek a szertartás megtartásával, még akkor sem, ha az pontosan el lett végezve. Néhány esetben a család részére az esemény körüli korlátozások miatt ajtóstul jön a stressz, ezt jogosak. - Ha a temetkezési vállalkozás többi rendezvényét szervezi, a temetkezés szervezőinek lehetősége nyílhat arra, hogy találkozzanak a temetőnél. Azonban ha valaki akarja, hogy ott legyenek családtagok vagy közel álló barátok is, tartsuk észben a helyzetet és az igényeket.
Table 32: one random ARC and MMLU translation examples each compared between OGX and OKAPI
Lang./Dataset Prompt
Score
English/MMLU When anatomically modern humans first arrived in the Middle East, who did they encounter? - Neandertals, the evolutionary descendants of the premodern human inhabitants of Europe and Asia - Homo erectus, the evolutionary descendants of the premodern human inhabitants of Europe and Asia - Homo sapiens, who had evolved independently in Asia and arrived there by 200,000 years ago - All of the above
German/OGX COMET: 0.8345 Auf wen trafen die anatomisch modernen Menschen bei ihrer ersten Ankunft im Nahen Osten? - Neandertaler, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo erectus, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo sapiens, der sich unabhängig von den Menschen in Asien entwickelt hatte und dort vor 200.000 Jahren ankam - Alle oben genannten
German/OKAPI COMET: 0.8402 Als anatomisch moderne Menschen erstmals im Nahen Osten ankamen, wen trafen sie? - Neandertaler, die evolutionären Nachkommen der frühzeitlichen menschlichen Bewohner Europas und Asiens - Homo erectus, die evolutionären Nachkommen der frühzeitlichen menschlichen Bewohner Europas und Asiens - Homo sapiens, die sich unabhängig in Asien weiterentwickelt hatten und bis vor 200.000 Jahren dort ankamen - Alle oben Genannten
English/ARC A class plans an investigation to see which brand of light bulb lasts the longest. Which of these steps should come first? - Repeat the investigation. - Write a report of the results. - Make a table for recording data. - Make daily observations of the light bulbs.
German/OGX COMET: 0.8546 Eine Klasse plant eine Untersuchung, um herauszufinden, welche Marke von Glühbirnen am längsten hält. Welcher dieser Schritte sollte zuerst erfolgen? - Wiederhole die Untersuchung. - Schreibe einen Bericht über die Ergebnisse. - Fertige eine Tabelle an, um die Daten festzuhalten. - Beobachte täglich die Glühbirnen.
German/OKAPI COMET: 0.8548 Eine Klasse plant eine Untersuchung, um herauszufinden, welche Marke von Glühbirnen am längsten hält. Welcher dieser Schritte sollte als Erstes durchgeführt werden? - Wiederholen Sie die Untersuchung. - Schreiben Sie einen Bericht über die Ergebnisse. - Erstellen Sie eine Tabelle zur Aufzeichnung von Daten. - Machen Sie tägliche Beobachtungen der Glühbirnen.

### v2

@@
-Towards Cross-Lingual LLM Evaluation for European Languages
+Towards Multilingual LLM Evaluation for European Languages
 Abstract
 The rise of Large Language Models (LLMs) has revolutionized natural language processing across numerous languages and tasks.
-However, evaluating LLM performance in a consistent and meaningful way across multiple European languages remains challenging, especially due to the scarcity of multilingual benchmarks.
-We introduce a cross-lingual evaluation approach tailored for European languages.
+However, evaluating LLM performance in a consistent and meaningful way across multiple European languages remains challenging, especially due to the scarcity of language-parallel multilingual benchmarks.
+We introduce a multilingual evaluation approach tailored for European languages.
 We employ translated versions of five widely-used benchmarks to assess the capabilities of 40 LLMs across 21 European languages.
@@
 The benchmarks and results are made publicly available to encourage further research in multilingual LLM evaluation.
-Towards Cross-Lingual LLM Evaluation for European Languages
-Klaudia Thellmann1 Bernhard Stadler1 Michael Fromm2,3 Jasper Schulze Buschhoff2 Alex Jude2 Fabio Barth4 Johannes Leveling2 Nicolas Flores-Herr2 Joachim Köhler2 René Jäkel1 Mehdi Ali2,3 1TU Dresden, Germany, 2 Fraunhofer IAIS, Germany 3Lamarr Institute, Germany, 4DFKI, Germany thanks: †Equal contribution.
+Towards Multilingual LLM Evaluation for European Languages
+Klaudia Thellmann1 Michael Fromm2,3 Bernhard Stadler1 Jasper Schulze Buschhoff2 Alex Jude2 Fabio Barth4 Johannes Leveling2 Nicolas Flores-Herr2 Joachim Köhler2 René Jäkel1 Mehdi Ali2,3 1TU Dresden, Germany, 2 Fraunhofer IAIS, Germany 3Lamarr Institute, Germany, 4 DFKI, Germany thanks: †Equal contribution.
 1 Introduction
-The development of Large Language Models (LLMs) has brought transformative advancements to natural language processing, enabling sophisticated tasks such as question-answering, summarization, and machine translation.
-However, despite these advances, the evaluation of LLMs across languages, especially those beyond English and Chinese, remains a significant challenge.
-Initiatives like CLEF [^1] https://clef2024.imag.fr/, FIRE [^2] http://fire.irsi.res.in/, and WMT [^3] https://www2.statmt.org/wmt24/ have focused on multilingual evaluations, yet there is a lack of dedicated benchmarks for systematically comparing LLM performance across the languages of Europa.
-Creating bespoke benchmarks for each language is a costly and time-consuming endeavor.
-This not only limits the scalability of evaluations but also results in a fragmented understanding of model performance across different languages.
-In many cases, evaluations are conducted in isolation, without a basis for meaningful comparison between languages.
-Moreover, human evaluation is widely recognized as the gold standard for assessing language model performance, as it captures nuances that automated benchmarks may miss.
-However, human evaluations are resource-intensive and difficult to scale across multiple languages.
-To address these challenges, we explore whether automatically translated benchmarks can provide a viable method to estimate human preferences in chat scenarios.
-While human evaluation would likely be the most accurate way to assess LLM performance, the resource demands of manual dataset creation and evaluation make it impractical for large-scale multilingual comparisons.
-Instead, this approach aims to reduce these resource demands while ensuring comprehensive evaluation coverage across all languages.
-In this paper, we focus on evaluating multilingual LLMs by utilizing translated versions of existing benchmarks.
-Our goal is to determine the efficacy of translated benchmarks and assess whether they can substitute manually generated ones, allowing for a more scalable and uniform evaluation across multiple languages.
+The development of Large Language Models (LLMs) has led to transformative advancements to natural language processing, enabling complex tasks such as question-answering, summarization, and machine translation.
+However, despite these advances, the evaluation of LLMs across languages—especially those beyond English and Chinese —remains a significant challenge.
+Initiatives like CLEF [^1] https://clef2024.imag.fr/, FIRE [^2] http://fire.irsi.res.in/, and WMT [^3] https://www2.statmt.org/wmt24/ have focused on multilingual evaluations, yet there is a lack of dedicated benchmarks that allow for consistent and comparable evaluation of LLM performance across different European languages.
+Creating custom benchmarks for each language is a costly and time-consuming, which limits the scalability of evaluations and can result in a fragmented understanding of model performance across different languages.
+Without comprehensive multilingual evaluations, comparisons between languages are often constrained, leading to an incomplete understanding of how models perform in languages beyond high-resource ones like English, German, or French.
+To address these challenges, we investigate whether automatically translated benchmarks can reliably assess model performance across languages and whether this performance correlates with human evaluations, similar to benchmarks conducted in English.
+In this paper, we focus on evaluating multilingual LLMs using translated versions of existing benchmarks.
+Our goal is to determine the effectiveness of these translated benchmarks and assess whether they can substitute manually generated ones, allowing for a more scalable and uniform evaluation across multiple languages.
 We make the following contributions:
   • We evaluate the performance of 40 state-of-the-art LLMs for five multilingual datasets covering English and 20 additional languages in Section 4.
-  • We assess whether translated benchmarks reflect human preferences to evaluate their viability in Section 5.
-  • We examine how the choice of translation service (i.e., DeeplPro and ChatGPT) affects the performance of current LLMs in accurately responding to various benchmark tasks in Section 6.
-  • We release EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K in 20 European languages as evaluation benchmarks for the NLP community [^4] https://huggingface.co/collections/openGPT-X/eu20-benchmarks-67093b13db8ff192dc39312d.
+  • We assess whether translated benchmarks correlate with human evaluations to asses their viability in Section 5.
+  • We compared translation services in Section 6, in particular DeeplPro and ChatGPT, to identify correlations in LLM performance and understand how translation quality impacts the models’ benchmark results.
+  • We release EU20-MMLU, EU20-HellaSwag, EU20-ARC, EU20-TruthfulQA, and EU20-GSM8K in 20 European languages as evaluation benchmarks for the NLP community.
+[^4] https://hf.co/collections/openGPT-X/eu20-benchmarks-67093b13db8ff192dc39312d.
 2 Related Work
-Multilingual benchmarks have predominantly been generated using two widely adopted methodologies, besides various intermediate approaches: either by using human annotators of multiple languages to develop a multilingual dataset from scratch Kocmi et al.
+Multilingual benchmarks have predominantly been generated using two widely adopted methodologies: either through human annotation to create datasets from scratch Kocmi et al.
 (2023); Goyal et al.
 (2022); Conneau et al.
-(2018) or by using LLMs to translate an existing benchmark into multiple languages Lai et al.
+(2018), or by translating existing benchmarks using large language models (LLMs) Lai et al.
 (2023); Tiedemann (2012).
-Creating a multilingual dataset with human annotators requires a significant amount of time, which impacts both the duration of production and the cost of the dataset Yang et al.
+While datasets developed by human annotators offer high-quality translations and task-specific accuracy, they require significant time and financial investment Yang et al.
 (2019).
-In addition, homogeneous benchmarks are required in each language instead of heterogeneous multilingual benchmarks for a cross-lingual analysis Lewkowycz et al.
-(2022); Tiedemann (2012).
-This method is increasingly inefficient because model scores can become saturated over time Ott et al.
-(2022), or benchmarks loose their relevance too fast if a desired performance is reached Lewkowycz et al.
+Furthermore, as models improve, these benchmarks can quickly become outdated, limiting their usefulness Ott et al.
+(2022); Srivastava et al.
 (2023).
-An alternative is the machine translation of existing benchmarks.
-This approach is both more cost-effective and time-efficient.
-However, the main disadvantage is the translation quality, which is highly dependent on the capability of the language model used Meng et al.
+In contrast, the machine translation of existing benchmarks is more efficient and cost-effective, but its quality is often inconsistent, especially for medium- and low-resource languages, where translation inaccuracies can impact model evaluation Meng et al.
+(2022); NLLB Team et al.
 (2022).
-In the past, some data sets have shown quality gaps in translation accuracy, with special consideration given to languages with medium and low resources Team et al.
-(2022).
-Those languages have only a few to very few data sets available for pre-training and evaluation.
-They are much more difficult to be translated by an LLM, when taking cultural nuances of each language in consideration.
-Translating these languages, in particular, can lead to imbalanced evaluations across the benchmarks Team et al.
-(2022); Goyal et al.
+These translation issues can lead to imbalanced evaluations across languages, particularly in languages with limited pre-training and evaluation resources Goyal et al.
 (2022); Conneau et al.
 (2018).
-Therefore, the selection of languages is a central decision to make when creating a multilingual benchmark.
+Thus, selecting the right languages and ensuring translation quality is critical for creating a robust multilingual benchmark.
 In addition to the choice of languages, it is crucial to consider the range of tasks that need to be covered for a benchmark.
-For example, creating a multilingual benchmark for a specific task provides valuable insights into model performance for that exact task but says little about a model’s general applicability Conneau et al.
+Creating a multilingual benchmark for a specific task provides valuable insights into model performance for that exact task but says little about a model’s general language capability Conneau et al.
 (2018).
-This also includes benchmarks that, for example, cover the question-answering (QA) task for a wide range of domains, as only the QA task is tested here Lewis et al.
+This also includes benchmarks that, for example, cover the question-answering (QA) task for a wide range of domains, as only the QA task is tested Lewis et al.
 (2020).
-Another approach is to cover multiple tasks to evaluate language models broadly.
-An example is the OKAPI dataset published by Lai et al.
-(Lai et al., 2023).
-For this benchmark Three datasets (AI2 Reasoning Challenge (ARC), HellaSwag, and MMLU) were translated into 26 languages: eight high-resource, 11 medium-resource, and seven low-resource languages Lai et al.
+Another important factor is the range of tasks covered in a benchmark.
+Task-specific benchmarks provide insights into a model’s performance for specific tasks but do not fully capture a model’s overall language capabilities Conneau et al.
+(2018); Lewis et al.
 (2020).
-Building on this approach, we have expanded the linguistic reach through additional translations with special consideration on European languages and increased the variety of tasks for a wider capability test of LLMs.
+To address this, benchmarks that span multiple tasks, such as the OKAPI dataset Lai et al.
+(2023), offer a broader evaluation of model capabilities.
+OKAPI, for example, translates four diverse tasks (ARC, HellaSwag, MMLU and TruthfulQA) into 31 languages, including high-, medium-, and low-resource languages.
+Building on this approach, we have expanded the linguistic reach through additional translations with special consideration on European languages and extend the variety of tasks.
+By including GSM8K Cobbe et al.
+(2021) we offer evaluations on multilingual grade school math word problems.
 Furthermore, we show that our data pre-processing has improved quality in certain language areas (See Section 3).
@@
 3.1 Translation Process
-We chose DeepL as translation service for its balance between translation accuracy and scalability.
+We chose DeepL as translation service due to its balance between translation accuracy and scalability.
 We translated five well-known datasets, ARC Clark et al.
@@
 (2020) from English into 20 European languages.
-These datasets encompass a mix of multiple-choice and open-ended generation tasks, each of which presented unique translation challenges.
-The translations maintained the original structure of each task to ensure the benchmarks were as consistent as possible across languages.
-Using DeepL’s XML-tag handling, we avoided issues with prompt formatting and ensured that key contextual elements, such as ’Question’ and ’Response’ prompts, were properly localized (cf.
+These datasets encompass a mix of multiple-choice and open-ended generation tasks, each presenting unique translation challenges.
+The translations preserved the original structure of each task to ensure consistency across languages.
+By leveraging DeepL’s XML tag handling, we avoided issues with prompt formatting and ensured that key contextual elements, such as ’Question’ and ’Response’ prompts, were appropriately localized (cf.
 Table 7).
-To enhance the visibility and accessibility of our benchmarks, we launched a multilingual leaderboard [^5] https://huggingface.co/spaces/openGPT-X/european-llm-leaderboard on Hugging Face, inspired by the popular Open LLM leaderboard [^6] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard.
+To enhance the visibility and accessibility of our benchmarks, we launched a multilingual leaderboard [^5] https://hf.co/spaces/openGPT-X/european-llm-leaderboard on Hugging Face, inspired by the popular Open LLM leaderboard [^6] https://hf.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard.
 3.2 Evaluation Process
-We conducted the model evaluations on the HPC Cluster Alpha Centauri (TUD/ZIH) [^7] https://compendium.hpc.tu-dresden.de, the JUWELS Booster (JSC) [^8] https://apps.fz-juelich.de/jsc/hps/juwels/index.html, and, for larger models, on the JURECA Cluster (WEST AI/JSC) [^9] https://westai.de/services/hardware/, using systems powered by NVIDIA A100 and H100 Tensor Core GPUs.
-For model evaluation, we used the LM Evaluation Harness Framework from Eleuther.AI [^10] https://github.com/OpenGPTX/lm-evaluation-harness in versions v0.4.1 and v0.4.3, evaluating in a multi-GPU setting using hf/accelerate with both data and model parallelism.
-For larger models, we deployed a local vllm-based inference server (v0.5.3) [^11] https://github.com/vllm-project/vllm/releases/tag/v0.5.5 using the LM Evaluation Harness local completions implementation (commit 42dc24/main) to perform pipeline-parallel evaluations across multiple GPUs on a single node.
-4 European LLM Evaluation Results
-This section provides an overview of the performance of large language models (LLMs) across various sizes, tasks and languages.
-Section 4.1 focuses on the evaluation across all 21 European languages and indicates which models provide a competitive performance in a multilingual setting.
-Section 4.2 shows the connection between the proportion of languages in the CommonCrawl dataset and the performance of the current best open weights models.
-Lastly, Section 4.3 highlights how the performance of the models depend on the language family.
-Given that many models provide little to no information about the composition of their training corpora, we focus our detailed analysis on the LLMs that demonstrate the most reliable performance across 21 languages.
-4.1 Global European LLM Evaluation
+We conducted the model evaluations on the HPC Cluster Alpha Centauri (TUD/ZIH) [^7] https://compendium.hpc.tu-dresden.de, the JUWELS Booster (JSC) [^8] https://apps.fz-juelich.de/jsc/hps/juwels/index.html, and, for larger models, on the JURECA Cluster [^9] https://westai.de/services/hardware/, using systems powered by NVIDIA A100 and H100 Tensor Core GPUs.
+For model evaluation, we used adapted versions of the LM Evaluation Harness Framework from Eleuther.AI in versions v0.4.1 and v0.4.3, evaluating in a multi-GPU setting using hf/accelerate with both data and model parallelism.
+For larger models, we deployed a local vLLM-based inference server (v0.5.3) [^10] https://github.com/vllm-project/vllm/releases/tag/v0.5.5 using the LM Evaluation Harness local-completions implementation111111commit 42dc24 LM Evaluation Harness to perform pipeline-parallel evaluations across multiple GPUs on a single node.
+We expended about 45,000 GPU hours conducting the model evaluations.
+4 Evaluation Results
+This section provides an overview of the performance of LLMs across various sizes, tasks and languages.
+Section 4.1 focuses on the evaluation of all 21 European languages considered in this paper and investigates which models provide a competitive performance in a multilingual setting.
+Section 4.2 explores the relationship between the proportion of languages in the CommonCrawl dataset and the performance of the best current open-weights model.
+Lastly, Section 4.3 examines how model performance varies across different language families.
+Since many models disclose little to no information about the composition of their training corpora, we focus our detailed analysis on the LLMs that exhibit the most consistent performance across the 21 languages.
+4.1 European LLM Evaluation
 This section provides an overview of the performance of several LLMs on various tasks and languages, grouped by their number of parameters in small-sized ($\leq$9B parameters), medium-sized ($>$9B & $<$27B parameters) and large ($>$27B parameters) models.
 The models are evaluated on our newly introduced tasks EU21-ARC, EU21-GSM8K, EU21-HellaSwag, EU21-MMLU, and EU21-TruthfulQA.
-An extensive overview of all models are provided in the Appendix as Tables (cf.
+A comprehensive overview of all models are provided in the Appendix, including Tables (cf.
 LABEL:tab:acc_std_21lang_1, LABEL:tab:acc_std_21lang_2) and task specific heatmaps (cf.
-Figure 5, 6 7, 8, 9).
+Figure 5,  6,  7,  8,  9).
 Figure 1: Accuracy results for four small sized LLMs on the EU21 tasks.
@@
 (2024), and Qwen2-7B-Instruct Yang et al.
-(2024), further models with a lower average performance like Gemma-7b Team et al.
-(2024a) are in the appendix LABEL:tab:acc_std_21lang_1 and LABEL:tab:acc_std_21lang_2
-Gemma-2-9b-Instruct consistently performs well (cf.
-Figure 1), by having a narrow interquartile ranges, signaling low variability and therefore stable performance in all languages.
-On the multilingual grade school math word benchmark EU20-GSM8K the median accuracy is worse than Meta-Llama-3.1-8B-Instruct, indicating lower math capabilities.
-In EU20-MMLU the model shows a robust behaviour, but the median is slightly lower than Qwen2-7B-Instruct.
-Both latter benchmarks indicate that the capacity of small models might not allow for both a reliable performance on all languages, aswell as the capability to handle specialized knowledge reliable.
-Meta-Llama-3.1-8B-Instruct also delivers solid results, occasionally surpassing Gemma-2-9b-Instruct, such as in the EU21-GSM8K task where it achieves the highest median performance.
-While capable in math related tasks, the model shows shortcomings across all 21 languages compared to Gemma-2-9b.
-Mistral-NeMo-Minitron-8B-Base performs competitively in certain tasks, particularly in more widely spoken languages, indicated by competitive maximum scores.
-However, its overall performance is weaker compared to other two models, with significant drops in less-resourced languages and complex tasks.
-In the EU21-GSM8K and EU21-MMLU tasks the median performance is the worst compared to the other models.
+(2024).
+Models with a lower average performance, like Gemma-7b Team et al.
+(2024a), are listed in Appendix Tables LABEL:tab:acc_std_21lang_1, LABEL:tab:acc_std_21lang_2.
+Gemma-2-9b-Instruct performs consistent across languages (cf.
+Figure 1), as indicated by its narrow interquartile range, but lags behind Meta-Llama-3.1-8B-Instruct on the EU21-GSM8K benchmark, signaling lower math abilities.
+In EU21-MMLU, the model shows a robust behaviour, but the median is slightly lower than Qwen2-7B-Instruct.
+The results on the latter two benchmarks indicate that the capacity of small models might not allow for a reliable performance on all languages and a good performance on specialized knowledge.
+Meta-Llama-3.1-8B-Instruct excels in math-related tasks, achieving the highest median performance in EU21-GSM8K, but its performance lags behind Gemma-2-9b-Instruct across all 21 languages.
+Mistral-NeMo-Minitron-8B-Base shows competitive performance in widely spoken languages but struggles in less-resourced languages and complex tasks.
+Its performance on EU21-GSM8K and EU21-MMLU is notably lower than the other models.
 Qwen2-7B-Instruct demonstrates the highest median accuracy in EU21-MMLU, but performs worse in EU21-HellaSwag and EU21-ARC.
@@
 This analysis (cf.
-Figure 2) focuses on medium-sized LLMs such as Mistral-Nemo-Instruct-12.2B [^12] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, Mixtral-8x7B-Instruct-v0.1 Jiang et al.
+Figure 2) focuses on medium-sized LLMs such as Mistral-Nemo-Instruct-12.2B [^12] https://hf.co/mistralai/Mistral-Nemo-Instruct-2407, Mixtral-8x7B-Instruct-v0.1 Jiang et al.
 (2024), Phi-3-medium-14B-4k-Instruct Abdin et al.
@@
 Mixtral-8x7B-Instruct-v0.1 shows the highest median performance on four out of the five tasks.
-Only on EU21-GSM8K its performance is lacking behind Mistral-Nemo-Instruct-12.2B, indicating worse performance on math related tasks.
-However, its performance tends to be more variable than Mistral-Nemo-Instruct-12.2B_2407, with some languages like Latvian displaying significantly lower scores, reflecting higher inconsistency across the board.
-Mistral-Nemo-Instruct-12.2B_2407 delivers solid results across all tasks and languages, with a particularly strong median performance in the EU21-GSM8K task.
-As evidenced by its narrow interquartile ranges the model signals low variability across all 21 languages compared to the other models.
-Phi-3-medium-14B-4k-Instruct presents competitive results mostly in English by having the highest GSM8K, MMLU values but lags vastly behind the previous two models in median peformance across all languages and tasks, indicating only a small amount of multilingual training corpora.
-Lastly, Vicuna-13-b-v1.5 based on the older LLama-2-13B is not competitive on any task, due to its missing multilingual capabilities.
-Compared to the small-sized category, Mixtral-8x7B-Instruct-v0.1 provides a substantial gain in EU21-Hellswag by achieving an median performance of 0.64, an absolute increase of 4% compared to Gemma-2-9b-Instruct.
+Only on EU21-GSM8K its performance is lacking behind Mistral-Nemo-Instruct-12.2B.
+However, its performance tends to be more variable across languages than Mistral-Nemo-Instruct-12.2B_240, indicated by the low minima in 4 out of 5 tasks.
+Mistral-Nemo-Instruct-12.2B_2407 delivers solid results across all tasks with a particularly strong median performance in the EU21-GSM8K task.
+As evidenced by its narrow interquartile ranges, the model signals low variability across all 21 languages compared to the other models.
+Phi-3-medium-14B-4k-Instruct presents competitive results in English by having the highest GSM8K and MMLU values, but lags vastly behind the previous two models in median peformance across all languages and tasks, indicating a focus on a few selected languages in the training corpora.
+Lastly, Vicuna-13-b-v1.5 based on the older LLama-2-13B is not competitive on any task, due to its limited two trillion English-centric training tokens.
+Compared to the small-sized category, Mixtral-8x7B-Instruct-v0.1 provides a substantial gain in EU21-HellaSwag by achieving an median performance of 0.64, an absolute increase of 4% compared to Gemma-2-9b-Instruct.
 On EU21-GSM8K Mistral-Nemo-Instruct-12.2B achieves 57% accuracy a small increase of 1% compared to Meta-Llama-3.1-8B-Instruct.
@@
 (2023), Meta-Llama-3.1-70B-Instruct Dubey et al.
-(2024) and c4ai-command-r-35B-v01131313https://huggingface.co/CohereForAI/c4ai-command-r-v01.
+(2024) and c4ai-command-r-35B-v01131313https://hf.co/CohereForAI/c4ai-command-r-v01.
 Gemma-2-27b-Instruct and Meta-Llama-3.1-70B-Instruct reliable outperforms the other models across most tasks, with median scores often exceeding an accuracy of 0.7.
 Their performance is not only high but also stable across different languages, as evidenced by its narrow interquartile ranges, signaling low variability.
-Probably due to the higher capacity (compared to Gemma-2-9b), Gemma-2-27b-Instruct can now also score higher on specialized knowledge tasks such as MMLU and GSM8K.
-Vicuna-33b-v1.3 and c4ai-command-r-35B-v01 perform worse compared to the previous two models. This indicates a monolingual training corpora paired with a monolingual tokenizer.
+We assume that due to the higher capacity (compared to Gemma-2-9b), Gemma-2-27b-Instruct now also scores higher on specialized knowledge tasks such as EU21-MMLU and EU21-GSM8K.
+Comparable, to the smaller Vicuna model also the 33 billion parameter version is not competitive on any task, due to its limited two trillion English-centric training tokens.
+c4ai-command-r-35B-v01 performs better in English compared to the Vicuna model.
 Especially Gemma-2-27b-Instruct convinces with a remarkable performance.
@@
 To better analyze the performance of LLMs across different languages, we categorized our target languages into high-resource (HRL) and medium-resource (MRL) groups, based on their proportion in the CommonCrawl dataset, which serves as the primary pre-training source for many LLMs.
-High-resource languages (HRL) had a CommonCrawl representation of 1% or more, whereas medium-resource languages (MRL) had between 0.01% and 1%.
+High-resource languages (HRL) have a CommonCrawl representation of 1% or more, whereas medium-resource languages (MRL) have between 0.01% and 1%.
 The language statistics accompanied with evaluation results of the LLama-3.1-70B-Instruct Dubey et al.
 (2024) model is depicted in Table 1.
-As can be seen there is a downwards trend with outliers in Swedish and Danish, both Germanic languages that might profit from English or other Germanic training corpora (cf. Table 1).
+There is a downwards trend with outliers in Swedish and Danish, both Germanic languages that might profit from English or other Germanic training corpora (cf.
+Table 1).
+The performance of EU21-TQA, EU21-MMLU, and EU21-GSM8K shows a slight decline in correlation with the number of speakers in millions and the percentage share of CommonCrawl data across different languages, while EU21-ARC and EU21-Hellaswag exhibit a more pronounced performance drop.
+Both EU21-ARC and EU21-Hellaswag are benchmarks that focus on reasoning, commonsense knowledge, and context understanding.
+These tasks might require more language specific knowledge, which may not be as well represented in languages with fewer speakers or less CommonCrawl data.
+In contrast, tasks like MMLU and GSM8K, while challenging, may rely more on factual recall, basic math, or language understanding.
+We hypothesize these are capabilities a model can easier transfer across languages.
+EU21- EU21- EU21- EU21- EU21-
 Lang. Speak. M [^14] https://www.ethnologue.com/insights/ethnologue200/ CC(%) [^15] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages Cat. Avg. ARC GSM8K HeSw MMLU TQA
@@
 Also Romanian as the smallest represented language achieves higher performance across all five tasks, indicating that the language family is of influence in the performance analysis.
-5 Correlation with Human Preference Scores
-We examine the correlation between our introduced EU21 benchmarks and the Elo scores from the LMSYS Chatbot Arena [^16] https://chat.lmsys.org/?leaderboard to better understand how human preferences are reflected.
-The Elo scores we use are derived from a dataset [^17] https://storage.googleapis.com/arena_external_data/public/clean_battle_20240814_public.json of about 1.8 million human quality judgements.
+5 Correlation with Human Preferences
+We examine the correlation between the introduced EU21 benchmarks and the Elo scores from the LMSYS Chatbot Arena [^16] https://chat.lmsys.org/?leaderboard to better understand how human preferences are reflected.
 In the LMSYS Chatbot Arena Chiang et al.
-(2024) users can vote on a pair of LLM responses in anonymized and randomized "battles".
-The battles are divided into several categories, some of which correspond to the response languages. The Elo rankings are computed independently for each category.
-In order to quantify how well the accuracies of a given model on our translated datasets correlate with the Elo scores, we compute the Pearson and Spearman Correlation Coefficients between the accuracy and the Elo scores of the $n=17$ different models.
-We compute these correlation separately for the Czech, French, German, Italian, Polish, Portuegese, Spanish and for English as a reference point.
-The remaining mostly lower ressource languages have a high uncertainty on their Elo scores as fewer battles (cf.
+(2024), users can vote on a pair of LLM responses in anonymized and randomized "battles".
+The resulting preference data are used to compute Elo rankings published on the LMSYS chatbot arena leaderboard.
+There are multiple independently computed rankings in different categories, some of which correspond to the languages of the conversations.
+We utilize an input dataset with timestamp 20240814 containing about 1.8 million human quality judgements provided for reproducing the rankings (cf.
+Section C.1).
+With these data, we compute the Elo scores using FastChat171717commit e5dc446f.
+The LLMs for the subsequent correlation analysis were mostly chosen using the following criteria:
+  • Text-only, Causal Transformer-based LM
+  • Checkpoints are publicly available on HF Hub
+  • Model Size below 100B parameters
+To quantify how well the accuracies of a given model on our translated datasets correlate with the Elo scores, Then we compute the Pearson and Spearman Correlation Coefficients between the accuracy and the Elo scores of $n=17$ models using scipy.stats with default settings and confidence level 95%.
+We compute these correlations for Czech, French, German, Italian, Polish, Portuguese, Spanish, as for these languages, at least 50% of the Elo ranks of the models we consider can be distinguished with about 95% confidence, and for English as a reference point.
+The remaining mostly smaller medium-resource languages have a high uncertainty in their Elo scores as fewer battles (cf.
 Table 21) are available.
-The LLMs for the analysis were mostly chosen based on the Elo rankings available, according to the following criteria:
-  • Text-only, Causal Transformer-based LM
-  • Checkpoints available on HF Hub
-  • Model Size below 100B parameters
-  • LMSYS Elo Ratings available for English, French, German and Spanish
 Lang. Pearson Spearman
@@
 (2023) and our evaluation datasets share a similar structure, they differ in key aspects such as translation methodology and sample alignment.
-The most notable difference between the two datasets is the translation approach: The OGX dataset was translated using DeepL, whereas OKAPI employed ChatGPT for this task.
+The most notable difference between the two datasets is the translation approach: Our dataset was translated using DeepL, whereas OKAPI employed ChatGPT for this task.
 The translations for the OKAPI tasks cover 31 languages, including both EU and non-EU languages.
-For the OpenGPT-X evaluation datasets, we focused on translations into 20 European languages.
-The common and new set of EU languages between the OKAPI and OpenGPT-X are shown in Table 3.
+For our evaluation datasets, we focused on translations into 20 European languages.
+The common and new set of EU languages between the OKAPI and our benchmarks are shown in Table 3.
 Common Languages New Languages
@@
 Swedish
-Table 3: Comparison of languages between OKAPI and OpenGPT-X Datasets
-In this section, we present a correlation analysis with a focus on the ARC, HellaSwag, and MMLU tasks, alongside a COMET score Rei et al.
+Table 3: Comparison of languages between OKAPI and our Datasets
+In this section, we present a correlation analysis with a focus on the EU20-ARC, EU20-HellaSwag, and EU20-MMLU tasks, alongside a COMET score Rei et al.
 (2023) evaluation to assess the translation quality across both datasets.
-The idea behind the correlation analysis is to record [^18] The evaluation was performed using the LM-Evaluation-Harness framework, with sample-level logging enabled.whether the model responded correctly or incorrectly for each OKAPI and OGX sample and to examine if the predictions are correlated, allowing direct performance comparison on sample-level on the same data.
-For a sample-level comparison, we ensured that OGX and OKAPI samples are matched and can be linked to their original translations.
-While ARC, HellaSwag, and MMLU samples have IDs matching the original English subsets, enabling mapping, this does not apply to TruthfulQA.
-For the correlation analysis, we calculated the correlation matrices for different models, tasks, and languages, listed in Appendix A.
+The idea behind the correlation analysis is to record [^18] The evaluation was performed using the LM-Evaluation-Harness framework, with sample-level logging enabled.whether the model responded correctly or incorrectly for each OKAPI sample and sample from our dataset and to examine if the predictions are correlated, allowing direct performance comparison on sample-level on the same data.
+For a sample-level comparison, we ensured that our examples and OKAPI samples are matched and can be linked to their original translations.
+While EU20-ARC, EU20-HellaSwag, and EU20-MMLU samples have IDs matching the original English subsets, enabling mapping, this does not apply to EU20-TruthfulQA.
+For the correlation analysis, we calculated the correlation matrices for different models, tasks, and languages, listed in Section D.3.
 The highest correlations, as shown in Table 30, are with Aya-23-8B (0.57), Meta-Llama-3-8B-Instruct (0.56), and gemma-7b (0.56), while bloom-7b1 (0.40) and bloomz-7b1 (0.44) show lower correlations.
 This indicates that Aya and Meta-Llama perform more consistently, while bloom-7b1 and bloomz-7b1 exhibit greater variability.
-As seen in Table 19, ARC has the highest correlation at 0.61, followed by Hellaswag and MMLU (both 0.54).
+As seen in Table 32, RC has the highest correlation at 0.61, followed by HellaSwag and MMLU (both 0.54).
 ARC’s structure and reliance on factual knowledge likely make model performance more stable and predictable.
-Regarding the task performance correlation, we can see from Table 18 that correlations range from 0.44 for Hungarian to 0.59 for Spanish.
-Hungarian’s lower score may reflect fewer training resources, while Spanish likely benefits from better data availability, leading to more consistent model performance.
+Table 31 shows that correlations range from 0.44 for Hungarian to 0.59 for Spanish.
+Hungarian’s lower score may reflect fewer training resources, while Spanish likely benefits from better data availability, leading to more consistent performance.
 Translation quality likely plays a role in these results, as ambiguous or incorrect translations, poorly adapted prompts, or task mismatches can negatively affect model performance.
 Thus, the observed correlations may also reflect the degree to which translation quality impacts each language’s evaluation results.
-Table 5, Table 6, Fig.
-4 and Table 7 The examples in Appendix A Table 5, Table 6, Fig.
-4 and Table 7 underscore the importance of context-aware translation in multilingual tasks, as naive translations can reduce clarity and coherence, ultimately affecting model performance.
-As part of the further analysis, we determined the contingency table (4) to compare the answers of the models on our evaluation datasets and the OKAPI datasets.
+Examples illustrating translation quality issues in the ten common languages can be found in LABEL:tab:EU21_okapi_hellaswag for the HellaSwag task.
+As part of the further analysis, we determined the contingency table (Table 4) to compare the answers of the models on our evaluation datasets and the OKAPI datasets.
 Both evaluations show similar alignment, with agreements on correct answers (38.1%) and incorrect answers (39.3%) being fairly balanced.
@@
 (2023) to analyze the translation quality of both datasets.
-Since there is no target translation for the different benchmarks, we evaluated the translation quality of the OGX and OKAPI datasets using the COMET-KIWI metric Rei et al.
+Since there is no target translation for the different benchmarks, we evaluated the translation quality of the our and OKAPI datasets using the COMET-KIWI metric Rei et al.
 (2023).
@@
 A score of 1 indicates a high-quality translation and a score close to 0 indicates a bad translation.
-Both the OGX and OKAPI datasets have a relatively high average score of 0.8208 and 0.8190, respectively, across all tasks and translations.
+Both our and OKAPI datasets have a relatively high average score of 0.8208 and 0.8190, respectively, across all tasks and translations.
 The exact scores can be found in the Appendix in Table Table 33.
-It is important to highlight that because there is no target translation, no statistical significance can be calculated between the scores of OKAPI and OGX.
-Therefore, we cannot conclude that the OGX translations are automatically higher quality based on the COMET KIWI score.
-Nevertheless, it can be shown that the OGX datasets have a consistently high translation quality across all languages (See Appendix in Table Table 29).
-We additionally compared all our models on the ARC benchmark Fig.
-10, Hellaswag Fig.
-11 and MMLU Fig.
-12.
-Overall we can see that all evaluated models achieve higher accuracies on our tasks compared to OKAPI.
-This analysis highlights the extent to which our and OKAPI results align across models, tasks, and languages, providing insights into the consistency and reliability of model evaluations across different datasets.
+It is important to highlight that because there is no target translation, no statistical significance can be calculated between the scores of OKAPI and ours.
+Therefore, we cannot conclude that our translations are automatically higher quality based on the COMET KIWI score.
+Nevertheless, it can be shown that our datasets have a consistently high translation quality across all languages (cf.
+Table Table 33).
+We additionally compared all our LLMs on the ARC, (Figure 10), HellaSwag (Figure 11) and MMLU benchmarks (Figure 12).
+Overall we can see that all evaluated LLMs achieve higher accuracies on our tasks compared to OKAPI.
+This analysis highlights the extent to which our and OKAPI results align across models, tasks, and languages, providing insights into the consistency and reliability of model evaluations across datasets.
 7 Conclusion
@@
 8 Limitations
-A hindrance in achieving benchmarks comparable across languages is localization of cultural aspects.
-Toxicity benchmarks in particular can suffer from this, as stereotypes and terms related to them likely do not translate directly been cultures and hence languages.
-To a lesser extent this applies to all benchmarks requiring some degree of common sense and world knowledge.
-A team of reseachers has launched an initiative [^19] https://huggingface.co/spaces/CohereForAI/MMLU-evaluationto correct some of these cultural biases in multilingual datasets (predominantly MMLU), through crowd-sourced manual work.
-While we are aware of this initiative, we opted not to conduct a similar enterprise in order to avoid doubled efforts.
-Aligning the results from the initative with our own translations will be an avenue for further work.
-Furthermore the degree to which a translation for an given task is not always feasible.
-The pronoun-resolution benchmark WinoGrande Sakaguchi et al.
-(2019) for instance is posed as a multiple choice task, where a on object or subject in a sentence is masked and each choice constitutes a replacement of the mask.
+It depends much on the task whether translations are a good way to create corpora in parallel languages.
+The ambiguity resolution benchmark WinoGrande Sakaguchi et al.
+(2019) for instance is posed as a multiple choice task, where a object or subject in a sentence is masked and each choice constitutes a replacement of the mask.
 The model is then tasked with selecting the sentence that is semantically correct.
-Tables LABEL:fig:winogrande_base and LABEL:fig:winogrande_spa demonstrate how in languages with gendered nouns or gendered adjectives in predicative use, the benchmark becomes easier due to strong grammatical signals indicating the correct choice of subject or object.
-It is for that reason we decided against translation of WinoGrande.
-In order to compare the translations of OGX with the original benchmark or other translations such as OKAPI, the samples must be mapped.
-This requires a one-to-one mapping.
-However, this mapping was impossible for all benchmarks; therefore, we could only compare three benchmarks with the OKAPI dataset: Hellaswag, MMLU, and ARC.
-For Hellaswag, we used the ind and the source_id; for ARC, we used the id; and for MMLU, the subtask_name and the natural sorting of the benchmarks, which turned out to be a one-to-one mapping.
-We analyzed the translated sequences with unexpectedly low scores (<0.5) during the quality analysis of the translations using the COMET KIWI score.
-Depending on the benchmark, some sequences were not translated.
-As a result, some answer options are either in the original language or do not appear at all.
-The maximum error rate is 2.3 % for Hellaswag, 1.8 % for ARC, and 2.1 % for MMLU across all languages.
-Although the errors show no significant difference in direct comparison with OKAPI (see Section 6) nor with the original benchmarks, the analysis using the COMET score in the translation pipeline should be considered in future work.
-No measures to indicate overfitting on English benchmarks, such as math-based benchmarks as in https://arxiv.org/pdf/2405.00332
+Table 5 and 6 in the Appendix demonstrate how in languages with gendered nouns or gendered adjectives in predicative use, the benchmark becomes easier due to strong grammatical signals indicating the correct choice of subject or object.
+Depending on the benchmark, some sequences were sporadically not translated, so some answer options are either in the original language (English) or are left blank.
+We detected the issue by analyzing the translated sequences with unexpectedly low COMET KIWI scores (<0.5) during the quality analysis of the translations.
+The maximum error rate for any language (assuming no errors for scores $\geq 0.5$) is 2.3% for HellaSwag (IT), 1.8% for ARC (ET), and 2.1% for MMLU (PT) across all languages.
+Variations in accuracy were noted when testing language models with different versions of the LM Eval Harness framework (v0.4.1, v0.4.3, and 42dc24/main).
+The differences (less than 1%), affected ARC and MMLU, were attributed mainly to changes in batch size, and task formatting.
+These variations emphasize the importance of maintaining consistent evaluation practices to ensure accurate model comparisons across different versions.
+Details can be found in section B.2.
+9 Ethical & Broader Impact
+The ability to evaluate large language models (LLMs) across a wide range of European languages, particularly underrepresented ones, is a critical step toward enhancing inclusivity and accessibility in natural language processing (NLP).
+By ensuring that LLMs can perform well in languages beyond English or other high-resource languages, we contribute to a more equitable digital landscape where speakers of less widely spoken languages have equal access to advanced language technologies.
+However, this inclusivity brings unique challenges, particularly in achieving benchmarks that are comparable across diverse linguistic and cultural contexts.
+Acknowledgments
+This work was funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine Learning and Artificial Intelligence, LAMARR22B as well as by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no.
+68GX21007D) and by the European Union’s Horizon 2020 research and innovation program under grant agreement No 101135671 (TrustLLM).
+The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V.
+(www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) as well as the Center for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for providing its facilities for automatic evaluation computations.
 Appendix A Experimental Setup
+A.1 Problematic Examples
+Tables 5 and 6 contain a sample from the development split of WinoGrande-XL and its translation, illustrating challenges in 1:1 translation of evaluation datasets.
+Sentence: The juice from the mango fruit could not fill up the cup because the _ is small.
+Choices:
+mango
+cup
+Table 5: Development split instance of the WinoGrande-XL dataset.
+Choices:
+El zumo de la fruta de mango no podía llenar la taza porque el mango es pequeño.
+El zumo de la fruta de mango no podía llenar la taza porque la taza es pequeña.
+Table 6: Potential Spanish translation of the instance in Table 5. Highlighted in red are differences between the two choices beyond replacement of the subject.
 Context: A cartoon animation video is shown with people wandering around and rockets being shot. two men
@@
 In contrast to the naive translation, this version properly adapts phrases, such as translating "rockets being shot" to "Raketen abgeschossen" (rockets launched), maintaining fluency and better capturing the meaning of the original text.
-Appendix B European LLM Evaluation Results
+Appendix B Evaluation Results
+B.1 Model overview
+Table 10: Model Overview; Legend: D: Used in downstream evaluation (cf. Section 4); T: Used in OKAPI translation comparison (cf. Section 6); C: Used in LMSYS correlation analysis (cf. Section 5)
+Model D T C Hugging Face Link (https://hf.co/...)
+Aya-23-8B CohereForAI/aya-23-8B
+Bloom-7b1 bigscience/bloom-7b1
+Bloomz-7b1 bigscience/bloomz-7b1
+c4ai-command-r-35B-v01 CohereForAI/c4ai-command-r-v01
+EuroLLM-1.7B utter-project/EuroLLM-1.7B
+EuroLLM-1.7B-Instruct utter-project/EuroLLM-1.7B-Instruct
+Gemma-7b google/gemma-7b-it
+Gemma-1.1-7b-Instruct google/gemma-1.1-7b-it
+Gemma-2-9b-Instruct google/gemma-2-9b-it
+Gemma-2-27b-Instruct google/gemma-2-27b-it
+Meta-Llama-2-7B meta-llama/Llama-2-7b-hf
+Meta-Llama-2-7B-Chat meta-llama/Llama-2-7b-chat-hf
+Meta-Llama-2-13B-Chat meta-llama/Llama-2-13b-chat-hf
+Meta-Llama-3-8B meta-llama/Meta-Llama-3-8B
+Meta-Llama-3-8B-Instruct meta-llama/Meta-Llama-3-8B-Instruct
+Meta-Llama-3.1-8B meta-llama/Llama-3.1-8B
+Meta-Llama-3.1-8B-Instruct meta-llama/Llama-3.1-8B-Instruct
+Meta-Llama-3.1-70B-Instruct meta-llama/Llama-3.1-70B-Instruct
+Mistral-7B-v0.1 mistralai/Mistral-7B-v0.3
+Mistral-7B-v0.3 mistralai/Mistral-7B-v0.3
+Mistral-7B-Instruct-v0.1 mistralai/Mistral-7B-Instruct-v0.1
+Mistral-7B-Instruct-v0.2 mistralai/Mistral-7B-Instruct-v0.2
+Mistral-7B-Instruct-v0.3 mistralai/Mistral-7B-Instruct-v0.3
+Mistral-Nemo-Base-12.2B_2407 mistralai/Mistral-Nemo-Instruct-2407
+Mistral-Nemo-Instruct-12.2B_2407 mistralai/Mistral-Nemo-Instruct-2407
+Mistral-NeMo-Minitron-8B-Base nvidia/Mistral-NeMo-Minitron-8B-Base
+Mixtral-8x7B-v0.1 mistralai/Mixtral-8x7B-v0.1
+Mixtral-8x7B-Instruct-v0.1 mistralai/Mixtral-8x7B-Instruct-v0.1
+Occiglot-7b-eu5 occiglot/occiglot-7b-eu5
+Occiglot-7b-eu5-Instruct occiglot/occiglot-7b-eu5-instruct
+Pharia-1-LLM-7B-control Aleph-Alpha/Pharia-1-LLM-7B-control-hf
+Pharia-1-LLM-7B-control-aligned Aleph-Alpha/Pharia-1-LLM-7B-control-aligned-hf
+Phi-3-medium-14B-4k-Instruct microsoft/Phi-3-medium-4k-instruct
+Phi-3-medium-14B-128k-Instruct microsoft/Phi-3-mini-128k-instruct
+Phi-3-mini-4k-Instruct microsoft/Phi-3-mini-4k-instruct
+Phi-3-mini-3.8B-128k-Instruct microsoft/Phi-3-mini-128k-instruct
+Qwen2-7B Qwen/Qwen2-7B
+Qwen2-7B-Instruct Qwen/Qwen2-7B-Instruct
+Vicuna-13b-v1.5 lmsys/vicuna-13b-v1.5
+Vicuna-33b-v1.3 lmsys/vicuna-33b-v1.3
+Table LABEL:tab:eval_model_overview contains an overview of the models used in the different evaluations performed in this paper.
+The columns D, T and C indicate whether the respective model was used in section 4, 6 and 5, respectively.
+The URLs for to the Hugging Face repositories of the models had to be shortened due to page width constraints and have to be prefixed with https://hf.co.
+B.2 Accuracy Variations
+Notable accuracy variations were observed when evaluating the performance of language models (LMs) using different versions of the LM Eval Harness (v0.4.1, v0.4.3, and 42dc24/main).
+The evaluations described in the following focused on the French version of the ARC-challenge dataset, and the findings highlight key differences and potential causes of these discrepancies.
+The results were conducted using Mistral-7B-v0.1, and accuracy was assessed across different batch sizes and floating-point formats.
+The model evaluations used batch sizes of 30, 32, and 60.
+These are the key observations and the numbers that underline the findings:
+Accuracy Variations Across Versions:
+Small but consistent discrepancies in accuracies were noted across the different versions.
+For instance, the accuracies in v0.4.1 and v0.4.3 for a batch size of 30 were both 40.27% for the ARC challenge, while the normalized accuracy varied slightly, i.e.
+v0.4.1 46.59% and v0.4.3 46.41%.
+When using a larger batch size of 60, both versions again reported an accuracy of 40.27%, with a normalized accuracy of v0.4.1 46.50% and v0.4.3 46.59%
+Base Version BS Acc_Norm Acc
+v0.4.1 30 0.46587 0.40273
+v0.4.1 32 0.46587 0.40188
+v0.4.1 60 0.46502 0.40273
+v0.4.3 30 0.46502 0.40273
+v0.4.3 32 0.46416 0.40188
+v0.4.3 60 0.46587 0.40273
+42dc24 30 0.46502 0.40273
+Table 11: Evaluation results with different LM Eval Harness base versions, batch sizes, and metrics.
+Potential Causes for Discrepancies:
+These differences suggest that changes in batch size and GPU counts influenced how padding samples were handled, causing minor variations in accuracy.
+This observation is consistent with the findings of other research groups using the LM Eval Harness for benchmarking  [^19] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard.
+Differences in the usage of random number generators in few-shot settings were noted, particularly between v0.4.1 and v0.4.3, which could affect model accuracy.
+In zero-shot settings, the accuracy difference was minimal, and evaluating in a few-shot setting did not significantly impact the outcome.
+Based on a manual code inspection, we suspect that few-shot results may vary depending on what other tasks are evaluated in the same run.
+Furthermore, sample logs revealed that in data-parallel mode, the same sequence of few-shot contexts is used in each process, such that only the actual test sample differed between the $i$-th samples of each process.
+This might possibly increase the standard error depending on how many processes (GPUs) are used as statistical noise caused by the specific selection of few-shot contexts might be amplified.
+We also compared float16 and float32 runs, and results showed no significant differences in accuracy between these formats, e.g.
+float16 accuracy 40.53% (in both versions) and float32 accuracy 40.44% (in both versions).
+Key Differences Identified:
+A primary source of accuracy discrepancies was found to be the changed reordering and collation of batches in v0.4.3.
+The accuracy results for reverse-patched reordering matched the v0.4.1 results, confirming that the reordering was the cause in this setting.
+Specifically, the v0.4.3 reverse-patched reordering exhibited 40.27% Accuracy and a Normalized Accuracy of 46.59% (exactly matching v0.4.1) Task specification fields previously treated as plain texts were re-interpreted as Jinja2 templates in v0.4.3, so trailing newlines that had been present in v0.4.1 were removed, leading to a slight accuracy shift in MMLU tasks.
+After patching, accuracy matched across both versions.
+Further Investigation:
+Few-shot settings showed matching results after reordering and newline patches were applied, ensuring consistency between v0.4.1 and v0.4.3 in few-shot evaluations.
+When multiplying the accuracies by the number of samples in the arc_challenge_FR test split (1172 samples), the higher accuracy resulted in 472 correct answers, while the lower accuracy yielded 471 correct answers.
+This marginal difference (0.085%) may be explained by numerical differences between log-probabilities during batching and ultimately the discontinuous nature of the Accuracy metric.
+Conclusion:
+The differences in batching and Jinja2 templating in v0.4.3 explain the accuracy variations observed in this experiment.
+The patched version of v0.4.3 aligns with v0.4.1 when these issues are addressed, suggesting that changes in batching or task description formatting may lead to non-deterministic results across different hardware or LM Eval Harness versions.
+While the accuracy variations were small (less than 1%), the discrepancies were linked to batching behavior and templating in v0.4.3.
+Ensuring consistency across these factors is crucial for accurate model evaluations, especially when comparing models across different versions.
+It should be a development goal for LLM evaluation frameworks that, except for padding, the inputs actually passed to the model under test should not vary depending on batch size, degree of data or model parallelism or other tasks evaluated at the same time.
+B.3 Downstream results tables
+Tables LABEL:tab:acc_std_21lang_1 and LABEL:tab:acc_std_21lang_2 contain model- and task-wise arithmetic means and population standard deviations across the EU21 languages, and the average of these means.
+Tables LABEL:tab:acc_std_lang_group_germanic_1, LABEL:tab:acc_std_lang_group_germanic_2, LABEL:tab:acc_std_lang_group_romance_1, LABEL:tab:acc_std_lang_group_romance_2, LABEL:tab:acc_std_lang_group_slavic_1, and LABEL:tab:acc_std_lang_group_slavic_2 contain the same information across the languages of the Germanic, Romance and Slavic language families instead of the EU21 languages.
+Table 20 lists the differences of the per-task mean accuracies of the high-resource languages minus the corresponding mean accuracies of the medium-resource languages according to table 1.
 Table 12: Performance of Models Across Tasks: Accuracy and Standard Deviation for 21 Languages
 Task Group
-Model Name Avg. ARC HellaSwag
+Model Name Avg. EU21-ARC EU21-HeSw
 Aya-23-8B Aryabumi et al. (2024) .441 .475 ± .134 .535 ± .145
@@
 Bloomz-7b1 Muennighoff et al. (2023) .287 .316 ± .101 .354 ± .115
+c4ai-command-r-35B-v01 .557 .592 ± .130 .646 ± .135
 EuroLLM-1.7B Martins et al. (2024) .338 .469 ± .036 .485 ± .038
@@
 Mistral-NeMo-Minitron-8B-Base Sreenivas et al. (2024) .502 .522 ± .114 .546 ± .120
-Mistral-Nemo-Base-12.2B_2407202020https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 .562 .615 ± .071 .641 ± .094
-Mistral-Nemo-Instruct-12.2B_2407212121https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 .598 .621 ± .071 .624 ± .093
+Mistral-Nemo-Base-12.2B_2407 .562 .615 ± .071 .641 ± .094
+Mistral-Nemo-Instruct-12.2B_2407 .598 .621 ± .071 .624 ± .093
 Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024) .591 .625 ± .138 .641 ± .137
 Mixtral-8x7B-v0.1 Jiang et al. (2024) .554 .608 ± .132 .636 ± .132
-Occiglot-7b-eu5222222https://huggingface.co/occiglot/occiglot-7b-eu5 .404 .470 ± .130 .511 ± .143
-Occiglot-7b-eu5-Instruct [^23] https://huggingface.co/occiglot/occiglot-7b-eu5-instruct .406 .484 ± .132 .519 ± .146
-Pharia-1-LLM-7B-control  [^24] https://huggingface.co/Aleph-Alpha/Pharia-1-LLM-7B-control .333 .393 ± .174 .433 ± .181
-Pharia-1-LLM-7B-control-aligned [^25] https://huggingface.co/Aleph-Alpha/Pharia-1-LLM-7B-control-aligned .346 .396 ± .177 .438 ± .185
+Occiglot-7b-eu5 .404 .470 ± .130 .511 ± .143
+Occiglot-7b-eu5-Instruct .406 .484 ± .132 .519 ± .146
+Pharia-1-LLM-7B-control .333 .393 ± .174 .433 ± .181
+Pharia-1-LLM-7B-control-aligned .346 .396 ± .177 .438 ± .185
 Phi-3-medium-14B-128k-Instruct Abdin et al. (2024) .508 .497 ± .168 .492 ± .166
@@
 Vicuna-33b-v1.3 Peng et al. (2023) .447 .527 ± .140 .547 ± .135
-c4ai-command-r-35B-v01262626https://huggingface.co/CohereForAI/c4ai-command-r-v01 .557 .592 ± .130 .646 ± .135
 Table 13: Performance of Models Across Tasks: Accuracy and Standard Deviation for 21 Languages
 Task Group
-Model Name MMLU TruthfulQA GSM8k
+Model Name EU21-MMLU EU21-TQA EU21-GSM8k
 Aya-23-8B .455 ± .070 .476 ± .028 .266 ± .105
@@
 Bloomz-7b1 .302 ± .039 .461 ± .037 .000 ± .001
+c4ai-command-r-35B-v01 .561 ± .074 .543 ± .032 .440 ± .085
 EuroLLM-1.7B .257 ± .006 .441 ± .027 .035 ± .006
@@
 Vicuna-33b-v1.3 .467 ± .078 .524 ± .039 .169 ± .082
-c4ai-command-r-35B-v01 .561 ± .074 .543 ± .032 .440 ± .085
 Table 14: Standard Deviation of Task Accuracies for Germanic Languages
 Task Group
-Model Name Avg. ARC HellaSwag
+Model Name Avg. EU21-ARC EU21-HeSw
 Aya-23-8B .509 .563 ± .105 .636 ± .121
@@
 Bloomz-7b1 .303 .344 ± .138 .379 ± .154
+c4ai-command-r-35B-v01 .627 .681 ± .076 .750 ± .087
 EuroLLM-1.7B .352 .507 ± .046 .518 ± .058
@@
 Vicuna-33b-v1.3 .532 .644 ± .072 .678 ± .103
-c4ai-command-r-35B-v01 .627 .681 ± .076 .750 ± .087
 Table 15: Standard Deviation of Task Accuracies for Germanic Languages
 Task Group
-Model Name MMLU TruthfulQA GSM8k
+Model Name EU21-MMLU EU21-TQA EU21-GSM8k
 Aya-23-8B .509 ± .050 .490 ± .026 .343 ± .077
@@
 Task Group
-Model Name Avg. ARC HellaSwag
+Model Name Avg. EU21-ARC EU21-HeSw
 Aya-23-8B .522 .600 ± .013 .668 ± .019
@@
 Bloomz-7b1 .337 .427 ± .106 .480 ± .130
+c4ai-command-r-35B-v01 .634 .712 ± .022 .766 ± .023
 EuroLLM-1.7B .346 .495 ± .019 .511 ± .024
@@
 Vicuna-33b-v1.3 .512 .627 ± .025 .641 ± .034
-c4ai-command-r-35B-v01 .634 .712 ± .022 .766 ± .023
 Table 17: Standard Deviation of Task Accuracies for Romance Languages
 Task Group
-Model Name MMLU TruthfulQA GSM8k
+Model Name EU21-MMLU EU21-TQA EU21-GSM8k
 Aya-23-8B .515 ± .007 .476 ± .029 .351 ± .026
@@
 Bloomz-7b1 .349 ± .035 .432 ± .040 .000 ± .001
+c4ai-command-r-35B-v01 .627 ± .009 .566 ± .026 .500 ± .023
 EuroLLM-1.7B .256 ± .008 .428 ± .030 .038 ± .006
@@
 Vicuna-33b-v1.3 .521 ± .010 .531 ± .012 .240 ± .024
-c4ai-command-r-35B-v01 .627 ± .009 .566 ± .026 .500 ± .023
 Table 18: Standard Deviation of Task Accuracies for Slavic Languages
 Task Group
-Model Name Avg. ARC HellaSwag
+Model Name Avg. EU21-ARC EU21-HeSw
 Aya-23-8B .412 .427 ± .097 .477 ± .105
@@
 Bloomz-7b1 .262 .259 ± .013 .294 ± .004
+c4ai-command-r-35B-v01 .535 .566 ± .079 .603 ± .087
 EuroLLM-1.7B .332 .455 ± .013 .472 ± .010
@@
 Vicuna-33b-v1.3 .443 .544 ± .058 .531 ± .050
-c4ai-command-r-35B-v01 .535 .566 ± .079 .603 ± .087
 Table 19: Standard Deviation of Task Accuracies for Slavic Languages
 Task Group
-Model Name MMLU TruthfulQA GSM8k
+Model Name EU21-MMLU EU21-TQA EU21-GSM8k
 Aya-23-8B .436 ± .047 .477 ± .030 .244 ± .056
@@
 Bloomz-7b1 .273 ± .016 .483 ± .014 .001 ± .000
+c4ai-command-r-35B-v01 .540 ± .045 .542 ± .021 .424 ± .053
 EuroLLM-1.7B .255 ± .005 .445 ± .026 .033 ± .004
@@
 Vicuna-33b-v1.3 .465 ± .019 .529 ± .047 .149 ± .033
-c4ai-command-r-35B-v01 .540 ± .045 .542 ± .021 .424 ± .053
-Model ARC GSM8K HellaSwag MMLU TruthfulQA
+Table 20: Differences between high and medium resource languages
+EU21 EU21- EU21- EU21- EU21-
+Model ARC GSM8k HeSw MMLU TQA
 Aya-23-8B .216 .155 .230 .109 .006
@@
 Bloomz-7b1 .128 -$0$141 .056 -0.025
+c4ai-command-r-35B-v01 .201 .124 .205 .112 .030
 EuroLLM-1.7B .052 .004 .050 -0.001 -0.031
@@
 Vicuna-33b-v1.3 .182 .117 .174 .101 .024
-c4ai-command-r-35B-v01 .201 .124 .205 .112 .030
-Table 16: Differences between high and medium resource languages
-Model Name PC
-Aya-23-8B 0.571
-Meta-Llama-3-8B-Instruct 0.562
-Mistral-7B-Instruct-v0.3 0.562
-Meta-Llama-3-8B 0.556
-gemma-7b 0.555
-Llama-2-7b-chat-hf 0.553
-Meta-Llama-3.1-8B 0.552
-gemma-1.1-7b-it 0.552
-Meta-Llama-3.1-8B-Instruct 0.549
-occiglot-7b-eu5 0.548
-occiglot-7b-eu5-instruct 0.546
-Mistral-7B-v0.1 0.542
-Mistral-7B-Instruct-v0.2 0.539
-Mistral-7B-v0.3 0.536
-Phi-3-mini-4k-instruct 0.534
-Mistral-7B-Instruct-v0.1 0.528
-Llama-2-7b-hf 0.520
-bloomz-7b1 0.439
-bloom-7b1 0.401
-Table 17: Model Performance Correlation Matrix (Sorted by Correlation Coefficient in Descending Order)
-(PC: Pearson Correleation Coefficient)
-Language PC
-es 0.589991
-fr 0.586389
-it 0.567793
-sv 0.564892
-da 0.557272
-nl 0.555682
-ro 0.546222
-de 0.545930
-sk 0.473632
-hu 0.444939
-Table 18: Language Correlation Coefficients (Sorted by Coefficient in Descending Order)
-(PC: Pearson Correleation Coefficient)
-Task PC
-EU21-ARC 0.610
-EU21-HeSw 0.543
-EU21-MMLU 0.538
-Table 19: Task Correlation Coefficients (Sorted by Coefficient in Descending Order)
-(PC: Pearson Correleation Coefficient)
+B.4 Downstream results heatmaps
 Figure 5: Task Accuracies per Language for EU21-ARC.
@@
 Figure 9: Task Accuracies per Language for EU21-GSM8k.
-Appendix C Correlation with Human Preference Scores
+Appendix C Correlation with Human Preferences
+C.1 Battle dataset overview
+The battle dataset [^20] https://storage.googleapis.com/arena_external_data/public/clean_battle_20240814_public.json (as of 2024-10-15) dataset we used for computing the Elo scores was published by LMSYS together with a Jupyter notebook [^21] https://colab.research.google.com/drive/1KdwokPjirkTmpO_P1WByFNFiqxWQquwH (as of 2024-10-15)for reproducibility purposes.
+Table 21 shows the proportions of the different languages in the dataset, according to the classification contained in the dataset itself:
 Language Number of Battles
@@
 Table 21: Battle Overview
-Model and Hugging Face Link
-c4ai-command-r-35b-v01
-gemma-1.1-7b-it
-gemma-2-27b-it
-gemma-2-9b-Instruct
-Llama-2-7b-chat-hf
-Llama-2-13b-chat-hf
-Meta-Llama-3-8B-Instruct
-Meta-Llama-3.1-8B-Instruct
-Meta-Llama-3.1-70B-Instruct
-Mistral-7B-Instruct-v0.1
-Mistral-7B-Instruct-v0.2
-Mixtral-8x7B-Instruct-v0.1
-Phi-3-medium-14b-4k-instruct
-Phi-3-mini-128k-Instruct
-Phi-3-mini-4k-Instruct
-Vicuna-13b-v1.5
-Vicuna-33b-v1.3
-Table 21: Model Overview
+C.2 Computation of model rank
+The model rank for a model $m_{A}$ is computed by the FastChat codebase as 1 plus the number of models $m_{B}$ such that the 97.5% quantile of the Elo score of $m_{A}$ is lower than the 2.5% quantile of the Elo score of $m_{B}$.
+For each such $m_{B}$, there is a 97.5% probability that the “true” Elo score of $m_{B}$ is greater than its 2.5% quantile, and this quantile is greater than the 97.5% quantile of the Elo score of $m_{A}$, which in turn has a 97.5% probability of being greater than the “true” Elo score of $m_{A}$.
+Thus, there is a probability greater than $97.5\%*97.5\%=95.0625\%$ that the “true” Elo score of $m_{B}$ is greater than the “true” Elo score of $m_{A}$.
+C.3 Elo Score Tables
+Tables LABEL:tab:lmsys_pearson_arc_few_shot, LABEL:tab:lmsys_pearson_gsm8k_few_shot, LABEL:tab:lmsys_pearson_hellaswag_few_shot, LABEL:tab:lmsys_pearson_mmlu_few_shot, and LABEL:tab:lmsys_pearson_truthfulqa_few_shot each contain the per-task language-wise correlations computed from the Elo scores along with p-values, confidence intervals, the number of models for which Elo scores were available, and the number of model ranks distinguishable with about 95% confidence.
+Tables LABEL:tab:lmsys_pearson_alltasks_few_shot and LABEL:tab:lmsys_pearson_alltasks_few_shot contains the same information for the average of the accuracy across all tasks, and LABEL:tab:lmsys_pearson_reasoning_few_shot for the average of the accuracy across all tasks except EU21-TQA.
 Table 22: Correlation Coefficients for Task EU20-ARC between Few-Shot Accuracy and LMSYS ELO Score, incl.
@@
 Appendix D Influence of translation service
+In this Section, we compare the OKAPI dataset with our dataset.
+We use the OKAPI datasets from the LM Eval Harness, which has been extended by seven languages compared to the original publication Lai et al.
+(2023).
+D.1 Dataset overview
+Table 29 compares the number of translated samples in the 11 languages that overlap with both datasets.
+Our dataset aims to cover all samples in different languages.
+For the validation splits, we cover all samples across all languages except for Hellaswag (99.93% translated).
+For the MMLU and ARC benchmark, the validation split is used for in-context examples in the LM Eval Harness.
+For the test split, we also covered all samples.
+The OKAPI translations cover, on average, 92.23% of the samples for all languages regardless of the split, excluding the MMLU validation split.
+For the MMLU validation split, 276 samples were translated on average.
+Table 29: Average sample number for each task of eleven translations.
+Split Original EU21 OKAPI
+MMLU Val. 1531 1531 276.18 ±$\pm 8$
+Test 14042 14042 13198.09 ±$\pm 101$
+Hellaswag Val 10042 10035.18 ±$\pm 19$ 9276.63 ±$\pm 112$
+ARC Val. 299 299 272.66 ±$\pm 0.5$
+Test 1172 1172 1071 ±$\pm 2$
+TruthfulQA Val. 817 817 782.09 ±$\pm 6$
+D.2 Accuracy Plots
+In Figures 10 , 11, and 12, we compare the accuracy of 19 LLMs on the EU21-ARC, EU21-Hellaswag, and EU21-MMLU with the OKAPI benchmarks, respectively.
+All models show a higher accuracy on our benchmarks regardless of the benchmark, with the only exception being the results of bloom-7b1 for the MMLU task (cf.
+Figure 12).
 Figure 10: Accuracy comparison of different models on our and OKAPIs translated ARC dataset.
@@
 Figure 12: Accuracy comparison of different models on our and OKAPIs translated MMLU dataset.
-Table 29: Comparison of ogx and okapi scores for ARC and Hellaswag benchmarks (Higher values in bold)
-Language ARC Hellaswag MMLU
-OGX OKAPI OGX OKAPI OGX OKAPI
+D.3 Correlations
+We have calculated the correlation using the Pearson correlation coefficient for a more detailed analysis of the results from Figures 10 , 11, and 12.
+In Table 30, 31, and 32 we show the model performance correlation matrix clustered by model name, language, and task, respectively.
+Table 30: Model Performance Correlation Matrix
+(PC: Pearson Correlation Coefficient)
+Model Name PC
+Aya-23-8B 0.571
+Meta-Llama-3-8B-Instruct 0.562
+Mistral-7B-Instruct-v0.3 0.562
+Meta-Llama-3-8B 0.556
+gemma-7b 0.555
+Llama-2-7b-chat-hf 0.553
+Meta-Llama-3.1-8B 0.552
+gemma-1.1-7b-it 0.552
+Meta-Llama-3.1-8B-Instruct 0.549
+occiglot-7b-eu5 0.548
+occiglot-7b-eu5-instruct 0.546
+Mistral-7B-v0.1 0.542
+Mistral-7B-Instruct-v0.2 0.539
+Mistral-7B-v0.3 0.536
+Phi-3-mini-4k-instruct 0.534
+Mistral-7B-Instruct-v0.1 0.528
+Llama-2-7b-hf 0.520
+bloomz-7b1 0.439
+bloom-7b1 0.401
+Table 31: Language Correlation Coefficients
+(PC: Pearson Correlation Coefficient)
+Language PC
+es 0.589991
+fr 0.586389
+it 0.567793
+sv 0.564892
+da 0.557272
+nl 0.555682
+ro 0.546222
+de 0.545930
+sk 0.473632
+hu 0.444939
+Table 32: Task Correlation Coefficients
+(PC: Pearson Correlation Coefficient)
+Task PC
+ARC 0.610
+HellaSwag 0.543
+MMLU 0.538
+D.4 Quality assurance
+In addition to the performance and correlation of the performance of the models, we also analyzed the translation quality using the COMET metric.
+Here, we performed the translations from our dataset and the OKAPI dataset for all available comparable languages.
+The metric we used was the COMET-KIWI metric, which is calculated using the “Unbabel/wmt22-cometkiwi-da” model [^22] https://huggingface.co/Unbabel/wmt22-cometkiwi-da.
+This model is reference-free and built on top of InfoXLM Rei et al.
+(2023).
+For Hellaswag and MMLU, the results of our translations are slightly higher than those of OKAPI with no evidence of significance.
+Table 33: Comparison of EU21 and Okapi scores for ARC and HellaSwag benchmarks (Higher values in bold)
+ARC Hellaswag MMLU
+Language EU21 OKAPI EU21 OKAPI EU21 OKAPI
 German 0.8289 0.8289 0.7857 0.7971 0.8110 0.8098
@@
 Czech 0.8456 - 0.7920 - 0.8178 -
-Table 30: Average sample number for each task of eleven translations.
-Split Original OGX OKAPI
-MMLU Val. 1531 276.18 ±$\pm 8$
-Test 14042 13198.09 ±$\pm 101$
-Hellaswag Val 10042 10035.18 ±$\pm 19$ 9276.63 ±$\pm 112$
-ARC Val. 299 299 272.66 ±$\pm 0.5$
-Test 1172 1172 1071 ±$\pm 2$
-TruthfulQA Val. 817 817 782.09 ±$\pm 6$
-Table 31: Hellaswag translation examples compared between OGX and OKAPI
+D.5 Translation examples
+The following are selected examples from the EU21 and OKAPI translations of the datasets.
+Table 34: Hellaswag translation examples compared between EU21 and OKAPI
 Lang./Dataset Prompt
@@
 [substeps] Check to see if the parent owned the budgie.
-German/OGX COMET: 0.6329 Wie man einen guten Wellensittich auswählt.
+German/EU21 COMET: 0.6329 Wie man einen guten Wellensittich auswählt.
 Suchen Sie nach Wellensittichen, die von einem seriösen Verkäufer verkauft werden.
@@
 - jumps into a black and white video and temporarily stops playing, saying, " big calibers ".
-French/OGX COMET: 0.5770 Une page de titre verte apparaît, avec des mots indiquant que la vidéo sera celle d’un joueur de saxophone ténor, ainsi que le site web permettant de recevoir des leçons.
+French/EU21 COMET: 0.5770 Une page de titre verte apparaît, avec des mots indiquant que la vidéo sera celle d’un joueur de saxophone ténor, ainsi que le site web permettant de recevoir des leçons.
 Un homme noir apparaît ensuite, avec un pull vert et un chapeau noir, jouant du saxophone.
@@
 - is drawn in with pick and shovel.
-Spanish/OGX COMET: 0.7803 Un hombre bebe de una lata junto a un agujero en la nieve.
+Spanish/EU21 COMET: 0.7803 Un hombre bebe de una lata junto a un agujero en la nieve.
 él - nada lentamente por una masa de agua poco profunda.
@@
 Moving around a bit can also minimize your nausea and help your stomach to settle.
-Italian/OGX COMET: 0.8592 Come alleviare un virus intestinale.
+Italian/EU21 COMET: 0.8592 Come alleviare un virus intestinale.
 Evitate i cibi solidi per qualche ora per permettere allo stomaco di calmarsi.
@@
 - seems very pleased with the information he has gathered as he allows the host to speak.
-Swedish/OGX COMET: 0.8174 En man med hörlurar befinner sig i en studio och talar i en mikrofon.
+Swedish/EU21 COMET: 0.8174 En man med hörlurar befinner sig i en studio och talar i en mikrofon.
 Andra män visas, varav en börjar spela munspel.
@@
 Invite the family to have other family members and close friends present if they wish, but be sensitive to their needs and to the space considerations.
-Hungarian/OGX COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni.
+Hungarian/EU21 COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni.
 Beszélj a családdal.
@@
 - lands on a pile of sand.
-Slovak/OGX COMET: 0.8336 Muž beží a pritom drží tyč na dráhe.
+Slovak/EU21 COMET: 0.8336 Muž beží a pritom drží tyč na dráhe.
 on - používa tyč na skok vo vzduchu.
@@
 - all clap and the girl smiles and turns toward the camera.
-Danish/OGX COMET: 0.8450 En lille pige løber væk fra pinataen, mens hun holder pinden.
+Danish/EU21 COMET: 0.8450 En lille pige løber væk fra pinataen, mens hun holder pinden.
 Barnet vender tilbage og rammer pinataen én gang.
@@
 Invite the family to have other family members and close friends present if they wish, but be sensitive to their needs and to the space considerations.
-Hungarian/OGX COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni.
+Hungarian/EU21 COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni.
 Beszélj a családdal.
@@
 Azonban ha valaki akarja, hogy ott legyenek családtagok vagy közel álló barátok is, tartsuk észben a helyzetet és az igényeket.
-Table 32: one random ARC and MMLU translation examples each compared between OGX and OKAPI
+Table 34: Hellaswag translation examples compared between EU21 and OKAPI
+Table 35: ARC and MMLU translation examples each compared between EU21 and OKAPI
 Lang./Dataset Prompt
@@
 - Neandertals, the evolutionary descendants of the premodern human inhabitants of Europe and Asia - Homo erectus, the evolutionary descendants of the premodern human inhabitants of Europe and Asia - Homo sapiens, who had evolved independently in Asia and arrived there by 200,000 years ago - All of the above
-German/OGX COMET: 0.8345 Auf wen trafen die anatomisch modernen Menschen bei ihrer ersten Ankunft im Nahen Osten?
+German/EU21 COMET: 0.8345 Auf wen trafen die anatomisch modernen Menschen bei ihrer ersten Ankunft im Nahen Osten?
 - Neandertaler, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo erectus, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo sapiens, der sich unabhängig von den Menschen in Asien entwickelt hatte und dort vor 200.000 Jahren ankam - Alle oben genannten
@@
 - Make daily observations of the light bulbs.
-German/OGX COMET: 0.8546 Eine Klasse plant eine Untersuchung, um herauszufinden, welche Marke von Glühbirnen am längsten hält.
+German/EU21 COMET: 0.8546 Eine Klasse plant eine Untersuchung, um herauszufinden, welche Marke von Glühbirnen am längsten hält.
 Welcher dieser Schritte sollte zuerst erfolgen?
@@
 - Machen Sie tägliche Beobachtungen der Glühbirnen.
+Table 35: one random ARC and MMLU translation examples each compared between EU21 and OKAPI
+Appendix E Other
+E.1 License
+MMLU, Hellaswag and GSM8K were released under MIT license; ARC under CC-BY-SA-4.0 license; TruthfulQA under Apache 2.0 license.
+EU20-MMLU, EU20-HellaSwag, EU20-GSM8K and EU20-TruthfulQA will be released with an permissive license, we ensure that all original licenses and notices are included.
+EU20-ARC will be released under the CC-BY-SA-4.0 license under the exactly the same terms as the original dataset ARC.
+E.2 Use of AI Assistants
+We acknowledge the use of ChatGPT-4 to generate code snippets for the Figures 1-3. The code was then extended and carefully reviewed.